### v3

# The canonical lamination calibrated by a cohomology class

Abstract.
Let $\rho$ be a unit cohomology class of degree $d-1$, on a closed oriented Riemannian manifold of dimension $d$. We construct a lamination $\lambda_{\rho}$ whose leaves are exactly the minimal hypersurfaces calibrated by every calibration in $\rho$. The geometry of $\lambda_{\rho}$ is closely related to the geometry of the unit ball of $H_{d-1}(M,\mathbb{R})$ when it is equipped with Gromov’s stable norm, so our main theorem constrains the shape of the stable unit ball in terms of the topology of $M$. These results establish a close analogy between the stable norm and Thurston’s earthquake norm on the tangent space to Teichmüller space.
Key words and phrases:
laminations, minimal hypersurfaces, calibrations, functions of least gradient, stable norm, Thurston asymmetric metric
2020 Mathematics Subject Classification:
primary: 49Q05; secondary: 53C38, 37F34
1. Introduction
Let $M$ be a closed oriented Riemannian manifold of dimension $2\leq d\leq 7$. The stable norm $\|\alpha\|_{1}$ of a homology class $\alpha\in H_{d-1}(M,\mathbf{R})$ is the infimum of the area of all $d-1$-cycles representing $\alpha$. The stable norm was introduced by Federer in his work [Federer1974], on the duality between area-minimizing currents and calibration cochains. Among other applications, the stable norm was studied by Gromov, [gromov2007metric], for its connections to systolic geometry and Brock and Dunfield, [Brock2017], because of its connection to the Thurston–Gromov simplicial norm.
A quarter-century ago, Auer and Bangert released a research announcement [Auer01], which proposed to study codimension-$1$ measured oriented laminations $\lambda$ in $M$ which minimize their mass in their homology class $[\lambda]\in H_{d-1}(M,\mathbf{R})$. [^1] We review the basic definitions related to laminations in §2.4. In our convention, all laminations are Lipschitz and nonempty. [^2] Some of Auer and Bangert’s work appears in an unpublished manuscript, [Auer12]. In codimension $1$, every homology class $\alpha$ can be represented by a mass-minimizing lamination $\lambda$ (whose mass then equals $\|\alpha\|_{1}$), which one can think of roughly think of as a canonical choice of representative of $\alpha$. If two laminations have common leaves, those leaves cannot intersect, and Auer and Bangert proposed to use this observation to establish a deep connection between the intersection theory of $M$ and the geometry of the stable unit ball. A similar approach was used by Balacheff and Massart, [Massart96, Massart2007], to study the stable unit ball when $M$ is a negatively curved surface.
While trying to prove the theorems claimed in [Auer01], it is often useful to imitate ideas of the works [Thurston98, Wolpert82, Gu_ritaud_2017, huang2024earthquakemetricteichmullerspace] of the Thurston school on Thurston’s asymmetric metric on Teichmüller space. To make this precise, let $g\geq 2$, let $\Sigma_{g}$ be the closed oriented surface of genus $g$, let $\mathscr{T}_{g}$ be its Teichmüller space, let $\rho,\sigma\in\mathscr{T}_{g}$ be hyperbolic metrics on $\Sigma_{g}$, and let $L(\rho,\sigma)\geq 1$ be the infimum of Lipschitz constants of maps $(\Sigma_{g},\rho)\to(\Sigma_{g},\sigma)$ homotopic to $\operatorname{id}_{\Sigma_{g}}$. Thurston’s stretch metric on $\mathscr{T}_{g}$ is $\log L$. Thurston’s stretch metric is studied using geodesic laminations on $(\Sigma_{g},\rho)$, and in particular the canonical maximally-stretched lamination given by the following theorem.
Theorem 1.1 ([Gu_ritaud_2017]).
For every $g\geq 2$ and $\rho,\sigma\in\mathscr{T}_{g}$, there exists a unique largest chain-recurrent geodesic lamination $\lambda_{\rho,\sigma}$ in $(\Sigma_{g},\rho)$ such that for every Lipschitz map $f:(\Sigma_{g},\rho)\to(\Sigma_{g},\sigma)$ homotopic to $\operatorname{id}_{\Sigma_{g}}$ such that $\mathrm{Lip}(f)=L(\rho,\sigma)$, $f$ stretches every leaf of $\lambda_{\rho,\sigma}$ by a factor of $L(\rho,\sigma)$.
The costable norm, $\|\cdot\|_{\infty}$, on $H^{d-1}(M,\mathbf{R})$ is the dual norm of the stable norm. The purpose of this paper is to flesh out the idea that the costable norm of $M$ is closely analogous to Thurston’s stretch distance in an infinitesimal neighborhood of a hyperbolic metric. Our main theorem is that for each class of unit costable norm, there is a lamination in $M$ of minimal hypersurfaces which is analogous to Thurston’s canonical lamination. Studying the structure of this lamination yields several of the proposed theorems of [Auer01]. A sample application of the theory we shall develop is that under purely topological assumptions, $M$ has many uniquely ergodic laminations of minimal hypersurfaces.
A calibration (of codimension $1$) is a closed $d-1$-form $F$ such that $\|F\|_{L^{\infty}}=1$. For any calibration $F$ on $M$, a hypersurface $N\subset M$ is $F$-calibrated if the pullback of $F$ to $N$ is the area form on $N$. If $N$ is $F$-calibrated, then the mean curvature of $N$ is 00, and if $N$ is closed then $N$ minimizes its area in its real homology class. This brings us to our main theorem:
Theorem 1.2.
For every $\rho\in H^{d-1}(M,\mathbf{R})$ such that $\|\rho\|_{\infty}=1$, there is a unique largest lamination $\lambda_{\rho}$ in $M$ such that for every calibration $F$ representing $\rho$, $F$ calibrates every leaf of $\lambda_{\rho}$.
The lamination $\lambda_{\rho}$ is the canonical lamination calibrated by $\rho$. We prove Theorem 1.2 in §4. The proof uses multiple results from [BackusCML], including the interpretation of mass-minimizing laminations in terms of functions of least gradient, functions $u$ which minimize their total variation $\int_{M}\star|\mathop{}\!\mathrm{d}u|$. We carefully note that the lamination $\lambda_{\rho}$ is not itself mass-minimizing, since it may not admit any sort of transverse measure and therefore does not have a well-defined homology class. However, any measured sublamination of $\lambda_{\rho}$ is mass-minimizing in its homology class.
To illustrate Theorem 1.2, suppose that $d=2$, so that we can identify homotopy classes of maps $M\to\mathbf{S}^{1}$ with homomorphisms $\pi_{1}(M)\to\mathbf{Z}$, which in turn can be identified with lattice points in $H^{1}(M,\mathbf{R})$. Let $\rho$ be such a lattice point; by rescaling $M$ we may assume that $\|\rho\|_{\infty}=1$. In that case, any calibration which represents $\rho$ is the derivative of a minimizing Lipschitz map in the homotopy class $\rho$, and every leaf of the canonical lamination calibrated by $\rho$ is maximally stretched by every minimizing Lipschitz map in $\rho$. Thus Theorem 1.2 is a generalization of a version of Theorem 1.1 where one works with homotopy classes $M\to\mathbf{S}^{1}$ rather than $[\operatorname{id}_{\Sigma_{g}}]$.
The definition of the costable norm makes sense in any cohomology group $H^{k}(M,\mathbf{R})$ and so it is natural to ask if Theorem 1.2 holds for $k\leq d-2$. If $k=1$, the analogue of Theorem 1.2 follows from Daskalopoulos and Uhlenbeck’s work [daskalopoulos2020transverse] on the $\infty$-Laplacian. However, if $22\leq k\leq d-$ 2, then the submanifolds calibrated by $\rho\in H^{k}(M,\mathbf{R})$ can intersect each other. For example, let $M=\mathbf{P}^{2}_{\mathbf{C}}$, let $\rho$ be the Kähler class of $\mathbf{P}^{2}_{\mathbf{C}}$, and let $F$ be the Kähler form of $\mathbf{P}^{2}_{\mathbf{C}}$, which is a calibration by Wirtinger’s inequality.
In §5, we study measured sublaminations of canonical calibrated laminations. This is motivated by the fact that the earthquake norm, the dual of the norm induced by Thurston’s stretch metric, is not strictly convex, and its failure to be strictly convex detects the failure of geodesic laminations to be uniquely ergodic [huang2024earthquakemetricteichmullerspace]. This suggests that if the stable norm is not strictly convex, then there should be canonical calibrated laminations which are not uniquely ergodic; it turns out that this is exactly what happens.
A finite Borel measure $\mu$ is transverse to a lamination $\lambda$, if $\operatorname{supp}\mu=\operatorname{supp}\lambda$ and $\mu$ is invariant under deformations which preserve the area forms of every leaf of $\lambda$. A transverse probability measure $\mu$ is ergodic if, for every Borel set $E$ which is a union of leaves of $\lambda$, either $\mu(E)=0$ or $\mu(E)=1$. The lamination $\lambda$ is uniquely ergodic, if there is a unique probability measure which is transverse to $\lambda$.
Let $\rho\in H^{d-1}(M,\mathbf{R})$ have unit norm. The canonical lamination $\lambda_{\rho}$ may not admit a transverse measure, but the proof of Theorem 1.2 shows that $\lambda_{\rho}$ has a sublamination which admits an ergodic transverse measure. Let
$B:=\{\alpha\in H_{d-1}(M,\mathbf{R}):\|\alpha\|_{1}\leq 1\}$
be the stable unit ball, and let
$\rho^{*}:=\{\alpha\in\partial B:\langle\rho,\alpha\rangle=1\}$
be the dual flat to $\rho$. Since the stable norm does not have to be convex, $\rho^{*}$ does not have to be a singleton.
Corollary 1.3.
For every $\rho\in H^{d-1}(M,\mathbf{R})$ with $\|\rho\|_{\infty}=1$, $\rho^{*}$ is the set of homology classes which are represented by probability measures which are transverse to sublaminations of the canonical lamination $\lambda_{\rho}$. Every extreme point of $\rho^{*}$ is represented by an ergodic measure on a sublamination of $\lambda_{\rho}$.
Auer and Bangert [Auer01] observed that one can use a lemma of Arnoux and Levitt [Arnoux1986] to estimate the number of ergodic measures on sublaminations of a lamination without closed leaves. So by Corollary 1.3, the Arnoux–Levitt lemma applied to $\lambda_{\rho}$ determines the structure of $\rho^{*}$. A homology class $\alpha\in H_{d-1}(M,\mathbf{R})$ has rational direction if there exists $c>0$ such that $c\alpha$ is in the image of the map $H_{d-1}(M,\mathbf{Z})\to H_{d-1}(M,\mathbf{R})$. Let $b_{1}(M):=\dim H_{1}(M,\mathbf{Q})$ be the first Betti number.
Theorem 1.4.
Let $F$ be a maximal flat of the stable unit sphere $\partial B$. Then:
  1. (1)
    $F$ is a convex polytope.
  2. (2)
    The number of vertices of $F$ with irrational direction is at most $\max(0,b_{1}(M)-1)$.
  3. (3)
    A vertex $\alpha$ of $F$ has rational direction iff $\alpha$ is represented by a closed leaf of $\lambda_{\rho}$.
For example, suppose that $M\cong\Sigma_{g}$ where $g\geq 2$. By a theorem of Massart [Massart1997StableNO], if a maximal flat $\rho^{*}$ has a point of rational direction, then $\rho^{*}$ has at most $33g-$ 3 vertices, all of which have rational direction. It follows that $\lambda_{\rho}$ consists of at most $33g-$ 3 closed geodesics, plus possibly a “spiraling” part which admits no transverse measures. On the other hand, if $\rho^{*}$ has no points of rational direction, then $\rho^{*}$ has at most $12g-$ 1 vertices, and $\lambda_{\rho}$ has no closed leaves.
Corollary 1.5.
If the stable unit ball $B$ is strictly convex, then every ergodic calibrated lamination is uniquely ergodic. In particular, if $b_{1}(M)\geq 2$ and $B$ is strictly convex, then all but countably many homology classes in $\partial B$ are represented by uniquely ergodic calibrated laminations without closed leaves.
The analysis of $\lambda_{\rho}$ yields the following theorem on the strict convexity of the stable unit ball, which was proposed without proof by Auer and Bangert [Auer01, Theorems 6 and 7]. The intersection product $\alpha\cdot\beta$ of two homology classes $\alpha,\beta$ is the Poincaré dual of the cup product $\operatorname{PD}(\alpha)\smile\operatorname{PD}(\beta)$, and the derived series of a group $\Gamma$ is defined by letting $\Gamma^{(0)}:=\Gamma$ and $\Gamma^{(n+1)}$ be the commutator subgroup of $\Gamma^{(n)}$.
Theorem 1.6.
One has:
  1. (1)
    If there is a line segment $[\alpha,\beta]\subset\partial B$, then $\alpha\cdot\beta=0$.
  2. (2)
    Let $\Gamma:=\pi_{1}(M)$. If $\Gamma^{(1)}/\Gamma^{(2)}$ is a torsion group, then $B$ is strictly convex.
For example, suppose that $M$ has the homotopy type of a torus. Then $B$ is strictly convex and so $M$ has many uniquely ergodic laminations of minimal hypersurfaces. If $M=\mathbf{R}^{d}/[0,1]^{d}$, then these laminations are the irrational foliations of $M$, but what Theorem 1.6 says is that the same behavior occurs regardless of the Riemannian metric on $M$.
Since the canonical calibrated lamination $\lambda_{\rho}$ does not have to be uniquely ergodic, it is natural to ask if there is a canonical measure on $\lambda_{\rho}$. In upcoming work [daskalopoulos2025], Daskalopoulos and Uhlenbeck will show there is a favored measure on any canonical maximally stretched lamination with only closed leaves. The same proof works on $\lambda_{\rho}$, but it is quite lengthy, so we omit the proof. For each $p<\infty$, there is a unique representative $F_{p}\in L^{p}(M,\Omega^{d-1})$ of $\rho$ which is $p$-harmonic; that is,
$\mathop{}\!\mathrm{d}F_{p}=0,\quad\mathop{}\!\mathrm{d}^{*}(|F_{p}|^{p-2}F_{p})=0.$
Theorem 1.7 ([daskalopoulos2025]).
Let $F_{p}$ be the $p$-harmonic representative of $\rho$, and let
$\mathop{}\!\mathrm{d}u_{p}:=|F_{p}|^{p-2}\star F_{p}.$
After taking a subsequence, $u_{p}$ converges in $L^{1}_{\mathrm{loc}}(\tilde{M})$ to a function $u$ of least gradient such that $\mu:=|\mathop{}\!\mathrm{d}u|\in\mathcal{M}(\lambda_{\rho})$. If $M$ is a hyperbolic surface and $\lambda_{\rho}$ only has closed leaves, then $\mu$ is independent of the choice of subsequence.
The proof of Theorem 1.7 suggests that if $\lambda_{\rho}$ has a uniquely ergodic sublamination $\kappa$ which has strictly larger Hausdorff dimension than the rest of $\lambda_{\rho}$ (and $M$ is not a closed hyperbolic surface), then $\mu$ conjecturally should be the unique measure on $\kappa$, and so should be independent of the choice of subsequence.
Theorems 1.2 and 1.7 are explicitly based on theorems about the earthquake norm, and there are also versions of Theorem 1.4 and 1.6 for the earthquake norm proven in [huang2024earthquakemetricteichmullerspace]. In §5.6 we conjecture a version of Corollary 1.3 for the earthquake norm.
Acknowledgments.
This work was closely inspired by ideas in [Auer01] of Franz Auer and Victor Bangert; I am especially grateful to Victor Bangert for allowing me to view their unpublished manuscript [Auer12]. I also thank Georgios Daskalopoulos and Karen Uhlenbeck for helpful discussions, and James Farre, Yi Huang, Zhenhua Liu, and Ben Lowe for helpful comments on an earlier draft. This research was supported by the National Science Foundation’s Graduate Research Fellowship Program under Grant No. DGE-2040433.
2. Preliminaries
2.1. Notation
Unless otherwise noted, $M$ always denotes a closed oriented Riemannian manifold of dimension $2\leq d\leq 7$. The operator $\star$ is the Hodge star on $M$. We denote the musical isomorphisms by $\sharp$,$\flat$ , ♭. We write $H^{\ell}$ for de Rham cohomology, but never a Sobolev space, which we instead denote $W^{\ell,p}$. The second fundamental form of a submanifold $N$ is $\mathrm{I\!I}_{N}$. If $K$ is a closed compact subset of a topological vector space, $\mathcal{E}(K)$ is the set of extreme points of $K$.
The sheaf of $\ell$-forms is denoted $\Omega^{\ell}$, and the sheaf of closed $\ell$-forms is denoted $\Omega^{\ell}_{\rm cl}$. We assume that $\ell$-forms are $L^{1}_{\mathrm{loc}}$, but not that they are continuous; hence $\mathop{}\!\mathrm{d}$ must be meant in the sense of distributions.
We write $A\lesssim_{\theta}B$ to mean that $A\leq CB$, where $C>0$ is a constant that only depends on $\theta$.
2.2. Differential forms in $L^{\infty}$
In this section, one can allow $M$ to be an arbitrary complete Riemannian manifold; compactness is unnecessary.
One of the main technical difficulties that we shall have to deal with is that we cannot prove the existence of continuous calibrations in general, and so we shall need to study differential forms which are merely in $L^{\infty}$. Such a form $F$ does not need to be well-defined on a set of zero measure, so in general, it does not make sense to integrate $F$ along a submanifold of $M$.
Theorem 2.1 ($L^{\infty}$ Poincaré lemma).
Let $x\in M$, and $10\leq k\leq d-$ 1. Then there exists $r_{*}>0$ which depends only on $\mathrm{Riem}_{M}$ near $x$ and the injectivity radius of $x$, such that for every $0<r\leq r_{*}$ and $F\in L^{\infty}(B(x,r),\Omega^{k+1}_{\rm cl})$, there exists a Hölder continuous $k$-form $A$ such that $F=\mathop{}\!\mathrm{d}A$.
Proof.
We may choose $r_{*}$ so that the exponential map $B_{\mathbf{R}^{d}}(0,r_{*})\to B(x,r_{*})$ is a diffeomorphism which induces topological isomorphisms for every function space under consideration. Thus it is no loss to replace $B(x,r)$ with the unit euclidean ball $\mathbf{B}^{d}$. By the main theorem of [Costabel2010], for every $1<p<\infty$ there is a continuous right inverse to the exterior derivative
[…]
The result now follows from the Sobolev embedding theorem if we take $p>d$. ∎
The next result is a rephrasing of [Anzellotti1983, Theorem 1.2], and asserts that closed $L^{\infty}$ $d-1$-forms can be integrated along Lipschitz hypersurfaces.
Theorem 2.2 (normal trace theorem).
Let $\iota:N\to M$ be the inclusion of an oriented Lipschitz hypersurface. Let $\mathcal{X}$ be the space of $F\in L^{\infty}(M,\Omega^{d-1})$ such that the components of $\mathop{}\!\mathrm{d}F$ are Radon measures. Then the pullback $\iota^{*}$ of $d-1$-forms extends to a bounded linear map
$\iota^{*}:\mathcal{X}\to L^{\infty}(N,\Omega^{d-1})$
satisfying the estimate
$\|\iota^{*}F\|_{L^{\infty}(N)}\leq\|F\|_{L^{\infty}(M)}.$ (2.1)
The comass of a differential $k$-form $F$ is
$\|F\|_{L^{\infty}_{*}}:=\sup_{\Sigma\subset M}\frac{1}{\mathrm{vol}(\Sigma)}\int_{\Sigma}F,$
where the supremum ranges over all oriented $k$-dimensional submanifolds $\Sigma$. It is clear that $\|F\|_{L^{\infty}_{*}}\leq\|F\|_{L^{\infty}}$, but if $F$ is a $d-1$-form, then the converse holds as well; we shall often use this fact without comment.
A $k$-current of finite mass is a continuous linear functional on the space
$C^{0}(M,\Omega^{d-k})\cap L^{\infty}(M,\Omega^{d-k})$
of bounded continuous $d-k$-forms. [^3] Be warned: this convention agrees with currents in algebraic geometry, where currents are viewed as generalizations of forms, but not geometric measure theory, where currents are viewed as generalizations of submanifolds. We denote the action of a current $T$ on a form $\varphi$ by $\int_{M}T\wedge\varphi$.
The mass of a $k$-current $T$ is
$\mathbf{M}(T):=\sup_{\|F\|_{L^{\infty}_{*}}\leq 1}\int_{M}T\wedge F.$
If $T$ represents a $d-k$-dimensional submanifold $\Sigma$, in the sense that $\int_{M}T\wedge F=\int_{\Sigma}F$, then $\mathbf{M}(T)=\mathrm{vol}(\Sigma)$. A function $u\in L^{1}_{\mathrm{loc}}(M)$ has bounded variation, denoted $u\in BV(M)$, if $\mathop{}\!\mathrm{d}u$ is a $1$-current of finite mass, in which case $\mathbf{M}(\mathop{}\!\mathrm{d}u)$ is the total variation of $u$, and we write $\int_{M}\star|\mathop{}\!\mathrm{d}u|$ to mean $\mathbf{M}(\mathop{}\!\mathrm{d}u)$.
One cannot multiply two arbitrary distributions, but one can define $\mathop{}\!\mathrm{d}u\wedge F$ when $u\in BV$, $F\in L^{\infty}$, and $\mathop{}\!\mathrm{d}F=0$. More precisely, we have:
Definition 2.3.
Let $u\in BV(M,\Omega^{k})$ and $F\in L^{\infty}(M,\Omega^{d-k-1})$. Assume that $\mathop{}\!\mathrm{d}F\in L^{d}(M,\Omega^{d-k})$. Then the Anzellotti wedge product of $\mathop{}\!\mathrm{d}u$ and $F$ is the distribution $\mathop{}\!\mathrm{d}u\wedge F$, such that for every test function $\chi\in C^{\infty}_{\mathrm{cpt}}(M,\mathbf{R})$,
$\langle\mathop{}\!\mathrm{d}u\wedge F,\chi\rangle:=-\int_{M}\chi u\wedge\mathop{}\!\mathrm{d}F-\int_{M}\mathop{}\!\mathrm{d}\chi\wedge u\wedge F.$
The next theorem is essentially [Anzellotti1983, Theorem 1.5], but we sketch the argument because Anzellotti did not formulate it in such generality.
Theorem 2.4 (Anzellotti’s theorem).
Let $u\in BV(M,\Omega^{k})$, $F\in L^{\infty}(M,\Omega^{d-k-1})$, and $\mathop{}\!\mathrm{d}F\in L^{d}(M,\Omega^{d-k})$. Then the Anzellotti wedge product $\mathop{}\!\mathrm{d}u\wedge F$ is well-defined as a distribution. In fact, $\mathop{}\!\mathrm{d}u\wedge F$ is a signed Radon measure, and
$\mathbf{M}(\mathop{}\!\mathrm{d}u\wedge F)\leq\mathbf{M}(\mathop{}\!\mathrm{d}u)\|F\|_{L^{\infty}_{*}}.$
In particular, if $k=0$,
$\mathbf{M}(\mathop{}\!\mathrm{d}u\wedge F)\leq\|F\|_{L^{\infty}}\int_{M}\star|\mathop{}\!\mathrm{d}u|.$ (2.2)
Proof.
By the $BV$ Sobolev embedding theorem, [evans2015measure, §5.6], for every $\chi\in C^{\infty}_{\mathrm{cpt}}(M)$, $\chi u$ belongs to the dual space of $L^{d}(M,\Omega^{d-k})$. Therefore for every $\chi\in C^{\infty}_{\mathrm{cpt}}(M)$, $\langle\mathop{}\!\mathrm{d}u\wedge F,\chi\rangle$ is finite, so $\mathop{}\!\mathrm{d}u\wedge F$ is well-defined as a distribution.
Suppose that $\operatorname{supp}\chi\Subset U$ for some $U\Subset M$. If $u$ is sufficiently smooth, then an integration by parts gives
$|\langle\mathop{}\!\mathrm{d}u\wedge F,\chi\rangle|=\left|\int_{M}\chi\mathop{}\!\mathrm{d}u\wedge F\right|\leq\|F\|_{L^{\infty}_{*}}\|\chi\|_{C^{0}}\mathbf{M}(1_{U}\mathop{}\!\mathrm{d}u).$
In general, we can find a sequence $(u_{n})\subset C^{\infty}$ such that $u_{n}\rightharpoonup^{*}u$ in $BV$. Then $u_{n}\rightharpoonup u$ in $L^{\frac{d}{d-1}}$ and $\mathop{}\!\mathrm{d}u_{n}\rightharpoonup^{*}\mathop{}\!\mathrm{d}u$ as currents of locally finite mass. Since we are testing $\mathop{}\!\mathrm{d}u$ against the $L^{d}$ form $\chi F$,
$|\langle\mathop{}\!\mathrm{d}u\wedge F,\chi\rangle|\leq\liminf_{n\to\infty}|\langle\mathop{}\!\mathrm{d}u_{n}\wedge F,\chi\rangle|\leq\|F\|_{L^{\infty}_{*}}\|\chi\|_{C^{0}}\liminf_{n\to\infty}\mathbf{M}(1_{U}\mathop{}\!\mathrm{d}u_{n}).$
But, by the portmanteau theorem [kechris2012classical, Theorem 17.20],
$\liminf_{n\to\infty}\mathbf{M}(1_{U}\mathop{}\!\mathrm{d}u_{n})\leq\liminf_{n\to\infty}\mathbf{M}(1_{\overline{U}}\mathop{}\!\mathrm{d}u_{n})\leq\mathbf{M}(\mathop{}\!\mathrm{d}u)$
which gives the desired estimate, since we only used the $C^{0}$ norm of $\chi$. ∎
2.3. Calibrated geometry
We recall calibrated geometry, which was developed by Harvey and Lawson [Harvey82].
Definition 2.5.
A calibration is a $k$-form $F$ such that $\mathop{}\!\mathrm{d}F=0$ and $\|F\|_{L^{\infty}_{*}}=1$. If $\Sigma$ is a $k$-dimensional submanifold, and $F$ pulls back to the Riemannian volume form of $\Sigma$, we say that $\Sigma$ is $F$-calibrated.
If $\Sigma$ is $F$-calibrated, then for any $k-1$-dimensional submanifold $\Lambda$,
$\mathrm{vol}(\Sigma)=\int_{\Sigma}F=\int_{\Sigma+\partial\Lambda}F\leq\mathrm{vol}(\Sigma+\partial\Lambda),$
so that $\Sigma$ is area-minimizing. On the other hand, if $A\in W^{1,\infty}(M,\Omega^{k-1})$, and $\Sigma$ is a closed $F$-calibrated submanifold, then
$\|F\|_{L^{\infty}_{*}}=1=\frac{1}{\mathrm{vol}(\Sigma)}\int_{\Sigma}F=\frac{1}{\mathrm{vol}(\Sigma)}\int_{\Sigma}F+\mathop{}\!\mathrm{d}A\leq\|F+\mathop{}\!\mathrm{d}A\|_{L^{\infty}_{*}},$
so $F$ minimizes its comass in its cohomology class if it calibrates a closed hypersurface.
The definition of $F$-calibration extends to currents. If $F$ is a calibration $k$-form, a $d-k$-current $T$ is $F$-calibrated if
$\int_{M}T\wedge F=\mathbf{M}(T).$
By Anzellotti’s theorem, Theorem 2.4, this definition makes sense as long as $T$ has locally finite mass. If $T$ is $F$-calibrated, then for any $d-k-1$-current $S$, $\mathbf{M}(T)\leq\mathbf{M}(T+\mathop{}\!\mathrm{d}S)$.
The comass and mass induce norms on cohomology and homology. The stable norm $\|\cdot\|_{1}$ on $H_{k}(M,\mathbf{R})$ is defined by
$\|\theta\|_{1}:=\inf_{\operatorname{PD}([T])=\theta}\mathbf{M}(T),$
where $\operatorname{PD}(\omega)$ is the Poincaré dual of the class $\omega$, and $T$ ranges over $d-k$-currents. The costable norm $\|\cdot\|_{\infty}$ is the dual norm of $\|\cdot\|_{1}$ on $H^{k}(M,\mathbf{R})$. The following theorem is a special case of the main theorem of [Federer1974, §4] but it is essential to us, so we sketch the proof.
Theorem 2.6.
For every $\rho\in H^{k}(M,\mathbf{R})$,
$\|\rho\|_{\infty}=\min_{[F]=\rho}\|F\|_{L^{\infty}_{*}},$
where $F$ ranges over closed measurable $k$-forms of class $\rho$.
Proof.
For each representative $F$ of $\rho$, and with $T$ ranging over all $d-k$-currents,
$\|\rho\|_{\infty}=\sup_{\begin{subarray}{c}\mathbf{M}(T)\leq 1\\ \mathop{}\!\mathrm{d}T=0\end{subarray}}\int_{M}T\wedge F,$
so in particular $\|\rho\|_{\infty}\leq\|F\|_{L^{\infty}_{*}}$. Conversely, for each $\kappa\in L^{1}(M,\Omega^{d-k})$ such that $\mathop{}\!\mathrm{d}\kappa=0$, let
$\Psi(\kappa):=\langle\rho,\operatorname{PD}([\kappa])\rangle.$
Then
$|\Psi(\kappa)|\leq\|\rho\|_{\infty}\|\operatorname{PD}([\kappa])\|_{1}\leq\|\rho\|_{\infty}\mathbf{M}(\kappa)$
so by the Hanh-Banach theorem, there exists $F\in L^{\infty}(M,\Omega^{k})$ such that $\|F\|_{L^{\infty}_{*}}\leq\|\rho\|_{\infty}$ and for every $\kappa\in L^{1}(M,\Omega^{d-k})$ such that $\mathop{}\!\mathrm{d}\kappa=0$,
$\int_{M}\kappa\wedge F=\langle\rho,\operatorname{PD}([\kappa])\rangle.$
This implies that $\mathop{}\!\mathrm{d}F=0$ and $[F]=\rho$. ∎
2.4. Laminations
We use roughly the same formalism for laminations as in [Morgan88], which we also used in [BackusCML]. Let $M$ be a Riemannian manifold. Fix an interval $I\subset\mathbf{R}$ and a box $J\subset\mathbf{R}^{d-1}$. A (codimension-$1$, Lipschitz) laminar flow box is a Lipschitz coordinate chart $\Psi:I\times J\to M$ and a compact set $K\subseteq I$, called the local leaf space, such that for each $k\in K$, $\Psi|_{\{k\}\times J}$ is a $C^{1}$ embedding, and the leaf $\Psi(\{k\}\times J)$ is a $C^{1}$ complete hypersurface in $\Psi(I\times J)$. Two laminar flow boxes belong to the same laminar atlas if the transition map preserves the local leaf spaces.
Definition 2.7.
A (codimension-$1$, Lipschitz) lamination $\lambda$ is a closed nonempty set $\operatorname{supp}\lambda$ and a maximal laminar atlas $\{(\Psi_{\alpha},K_{\alpha}):\alpha\in A\}$ such that
$\operatorname{supp}\lambda\cap\Psi_{\alpha}(I\times J)=\Psi_{\alpha}(K_{\alpha}\times J).$
Note carefully that the leaves of a lamination will typically not be embedded, but merely injectively immersed. The following theorem allows us to construct laminations without explicitly constructing their flow boxes, provided that the leaves are minimal hypersurfaces.
Theorem 2.8 ([BackusCML, Theorem A]).
Let $\mathcal{S}$ be a set of disjoint minimal hypersurfaces in $M$, such that $\bigcup_{N\in\mathcal{S}}N$ is a closed set, and $\sup_{N\in\mathcal{S}}\|\mathrm{I\!I}_{N}\|_{C^{0}}<\infty$. Then $\mathcal{S}$ is the set of leaves of a Lipschitz lamination $\lambda$, such that the normal vector to the leaves of $\lambda$ extends to a Lipschitz section of a line bundle on $M$.
In our application, $d\leq 7$ and the hypersurfaces in $\mathcal{S}$ are stable, so we can check the hypothesis on curvature in Theorem 2.8 using [Schoen81]. In my experience, it is a common misconception that the hypothesis on curvature can be removed, but the next example shows that it cannot be.
Example 2.9.
Let $M$ be the unit ball of $\mathbf{R}^{3}$, let
$\iota_{n}:\mathbf{R}\times\mathbf{S}^{1}$ $\to\mathbf{R}^{3}$
$(z,\theta)$ $\mapsto(2^{-n}\cosh(2^{n}z)\cos\theta,2^{-n}\cosh(2^{n}z)\sin\theta,z+1/n)$
and let $N_{n}:=M\cap(\iota_{n})_{*}(\mathbf{R}\times\mathbf{S}^{1})$. So if $\iota_{n}(z,\theta)\in N_{n}$ then $2^{-n}\cosh(2^{n}z)\leq 1$. In this case, if $n\geq 3$ then
$2^{n}z\leq\operatorname{arcosh}(2^{n})\leq n,$
and in particular
$N_{n}\subset\{(x,y,z)\in M:1/n-n/2^{n}\leq z\leq 1/n+n/2^{n}\}.$
It follows that if $n,m\geq 12$ then $N_{n}\cap N_{m}=\emptyset$. Let $\mathcal{S}$ be the set of all $N_{n}$s where $n\geq 12$, and the $z$-axis; one could call such a structure a stack of catenoids. A stack of catenoids is not a lamination, even though its leaves are disjoint and minimal, and have closed union. The leaves of a stack of catenoids have Morse index $\leq 1$ and area $\leq 5\pi$, which is “almost as good” as having bounded Gaussian curvature, but even this is not enough.
An arbitrary lamination cannot be viewed as a current, but following Ruelle and Sullivan [Ruelle75], we view laminations which have been equipped with transverse measures and orientations as currents, so our next task is to define the Ruelle-Sullivan current.
Definition 2.10.
Let $\lambda$ be a lamination with laminar atlas $\{(\Psi_{\alpha},K_{\alpha}):\alpha\in A\}$. Then:
  1. (1)
    $\lambda$ is equipped with an orientation if the transition maps $\Psi_{\alpha}^{-1}\circ\Psi_{\beta}$ are orientation-preserving.
  2. (2)
    A transverse measure $\mu$ to $\lambda$ consists of Radon measures $\mu_{\alpha}$ on each local leaf space $K_{\alpha}$, such that the transition maps $\Psi_{\alpha}^{-1}\circ\Psi_{\beta}$ send $\mu_{\beta}$ to $\mu_{\alpha}$, and $\operatorname{supp}\mu_{\alpha}=K_{\alpha}$. The pair $(\lambda,\mu)$ is a measured lamination.
  3. (3)
    Suppose that $\lambda$ is oriented, $\mu$ is a transverse measure to $\lambda$, and $\{\chi_{\alpha}:\alpha\in A\}$ is a partition of unity subordinate to $\{\Psi_{\alpha}(I\times J):\alpha\in A\}$. The Ruelle-Sullivan current $T_{\mu}$ acts on $\varphi\in C^{0}_{\mathrm{cpt}}(M,\Omega^{d-1})$ by
    $\int_{M}T_{\mu}\wedge\varphi:=\sum_{\alpha\in A}\int_{K_{\alpha}}\left[\int_{\{k\}\times J}(\Psi_{\alpha}^{-1})^{*}(\chi_{\alpha}\varphi)\right]\mathop{}\!\mathrm{d}\mu_{\alpha}(k).$
It bears repeating that in our convention, a lamination $\lambda$ is nonempty, and if $\mu$ is a transverse measure to $\lambda$, then $\operatorname{supp}\mu=\operatorname{supp}\lambda$.
Let $(\lambda,\mu)$ be a measured oriented lamination. It is a straightforward modification of the arguments of [daskalopoulos2020transverse, §8] to show that the Ruelle-Sullivan current $T_{\mu}$ is a closed $1$-current which is well-defined, in the sense that $T_{\mu}$ does not depend on the choice of laminar atlas. Furthermore, by [BackusCML, Lemma 3.1],
$T_{\mu}=\mathbf{n}_{\lambda}^{\flat}\mu$ (2.3)
where $\mathbf{n}_{\lambda}^{\flat}$ is the conormal $1$-form to $\lambda$ and $\mu(U):=\int_{U}\star|T_{\mu}|$ for every open set $U$. Often we leave $\mu$ implicit and just write $T_{\lambda}$ for $T_{\mu}$.
Let $\lambda$ be a lamination. A Borel set $E\subseteq\operatorname{supp}\lambda$ is saturated if, for every leaf $N$ of $\lambda$ such that $N\cap E$ is nonempty, $N\subseteq E$. Every leaf of $\lambda$ is Borel, and therefore saturated. A sublamination of $\lambda$ is a closed saturated set. Every sublamination of $\lambda$ the flow boxes of $\lambda$ and therefore is itself a lamination.
Lemma 2.11.
Let $\mathscr{S}$ be a nonempty set of laminations. Suppose that there exists a hypersurface which is a leaf of every lamination in $\mathscr{S}$. Then there exists a lamination whose set of leaves is the intersection of the sets of leaves of the laminations in $\mathscr{S}$.
Proof.
Let $\lambda\in\mathscr{S}$, and let $(\Psi_{\alpha},K_{\alpha})_{\alpha\in A}$ be a laminar atlas for $\lambda$. Let $K^{\prime}_{\alpha}$ be the set of $k\in K_{\alpha}$ such that for every $\kappa\in\mathscr{S}$, there exists a leaf $N$ of $\kappa$ such that
$(\Psi_{\alpha})_{*}(\{k\}\times J)\subseteq N.$
It is clear that this property is preserved by transition maps. Then $K_{\alpha}^{\prime}$ is an intersection of compact sets (since the local leaf space of each $\kappa\in\mathscr{S}$ is compact), so $K_{\alpha}^{\prime}$ is compact. The hypersurface which is a common leaf of every lamination in $\mathscr{S}$ witnesses that for some $\alpha$, $K_{\alpha}^{\prime}$ is nonempty. Therefore $(\Psi_{\alpha},K^{\prime}_{\alpha})_{\alpha\in A}$ is a laminar atlas. The fact that $K_{\alpha}^{\prime}$ is compact for every $\alpha$ implies that the supposed lamination whose atlas is $(\Psi_{\alpha},K^{\prime}_{\alpha})_{\alpha\in A}$ has a closed support. ∎
We shall also need a form of the Morgan–Shelan decomposition, [Morgan88, Theorem I.3.2], of a measured lamination. To formulate it, let us say that a lamination $\lambda$ is exceptional [^4] Exceptional laminations are often called minimal, but that clashes with the use of the word “minimal” to refer to vanishing mean curvature, so we have not adopted this terminology. if every leaf of $\lambda$ is dense in $\operatorname{supp}\lambda$, and $\lambda$ is not a single closed leaf. A lamination $\lambda$ is a parallel family of closed leaves if there exists a closed leaf $N$ of $\lambda$ with trivial normal bundle, such that every leaf of $\lambda$ is a section of the normal bundle of $N$.
Theorem 2.12 (Morgan–Shelan decomposition).
Suppose that $M$ is closed and oriented. For every measured oriented lamination $\lambda$, one of the following holds:
  1. (1)
    $\lambda$ is a foliation with a dense leaf.
  2. (2)
    $\lambda$ is the disjoint union of finitely many clopen sublaminations $\kappa$, such that either $\kappa$ is exceptional, or $\kappa$ is a parallel family of closed leaves.
Proof.
First observe that the proof of [Morgan88, Theorem I.3.2] goes through for any lamination $\lambda$ such that no leaf of $\lambda$ is dense in $M$, even if $\lambda$ is a foliation. It then remains to rule out the case that $\kappa$ is a family of sections of a nontrivial normal bundle of a closed leaf: this holds because $\lambda$ is oriented. ∎
3. Calibrated laminations and functions of least gradient
3.1. Calibrated laminations
Let $M$ be a closed oriented Riemannian manifold. Let $F$ be a calibration on $M$, and $\lambda$ a measured oriented lamination in $M$. There are two things that one could conceivably mean by saying that $\lambda$ is $F$-calibrated: that every leaf of $\lambda$ is $F$-calibrated, or that the Ruelle-Sullivan current, $T_{\lambda}$, is $F$-calibrated. The purpose of this section is to show that these two notions are equivalent.
Definition 3.1.
Let $F\in L^{\infty}(M,\Omega^{d-1})$ be a calibration. A lamination $\lambda$ is $F$-calibrated if every leaf of $\lambda$ is $F$-calibrated.
By the normal trace theorem, Theorem 2.2, this definition makes sense. Of course, one is only really interested in calibrated laminations if they are mass-minimizing, so now we recall that the mass of a measured oriented lamination $\lambda$ is
$\mathbf{M}(\lambda):=\mathbf{M}(T_{\lambda}).$
Since a current can be approximated by smooth $1$-forms in the weakstar topology on currents, every current has a cohomology class $[T]\in H^{1}(M,\mathbf{R})$. Thus, the homology class $[\lambda]\in H_{d-1}(M,\mathbf{R})$ is the Poincaré dual of $[T_{\lambda}]$.
Definition 3.2.
Let $\lambda$ be a measured oriented lamination, and assume that $M$ is compact. Then $\lambda$ is homologically minimizing, if for every measured oriented lamination $\kappa$ such that $[\lambda]=[\kappa]$,
$\mathbf{M}(\lambda)\leq\mathbf{M}(\kappa).$
Let $(\lambda,\mu)$ be a measured oriented lamination. Let $(\chi_{\alpha})$ be a locally finite partition of unity subordinate to a laminar atlas $(U_{\alpha},K_{\alpha})$ for $\lambda$. If $\sigma_{\alpha,k}$ denotes the leaf in $U_{\alpha}$ corresponding to the real number $k\in K_{\alpha}$, then the definition of the Ruelle-Sullivan current unpacks as
$\int_{M}T_{\lambda}\wedge F=\sum_{\alpha}\int_{K_{\alpha}}\int_{\sigma_{\alpha,k}}\chi_{\alpha}F\mathop{}\!\mathrm{d}\mu_{\alpha}(k).$ (3.1)
Since $T_{\lambda}$ and $F$ are closed, if $M$ is closed, then the left-hand side of (3.1) is a homological invariant:
$\int_{M}T_{\lambda}\wedge F=\langle[F],[\lambda]\rangle.$ (3.2)
Lemma 3.3.
Let $F$ be a calibration. Let $T_{\lambda}$ be the Ruelle-Sullivan current of a measured oriented lamination $\lambda$. Then the following are equivalent:
  1. (1)
    $T_{\lambda}$ is $F$-calibrated.
  2. (2)
    $\lambda$ is $F$-calibrated.
Proof.
First suppose that $T_{\lambda}$ is $F$-calibrated. Let $(\chi_{\alpha})$ be a locally finite partition of unity subordinate to an open cover $(U_{\alpha})$ of flow boxes for $\lambda$, let $(K_{\alpha})$ be the local leaf spaces, and let $(\mu_{\alpha})$ be the transverse measure. After refining $(U_{\alpha})$ we may assume that $U_{\alpha}$ is a ball which satisfies the hypotheses of the $L^{\infty}$ Poincaré lemma, Theorem 2.1. After shrinking $U_{\alpha}$ we may assume that $\chi_{\alpha}>0$ on $U_{\alpha}$. Then for leaves $\sigma_{\alpha,k}$, we rewrite (3.1) as
$\mathbf{M}(\lambda)=\int_{M}T_{\lambda}\wedge F=\sum_{\alpha}\int_{K_{\alpha}}\int_{\sigma_{\alpha,k}}\chi_{\alpha}F\mathop{}\!\mathrm{d}\mu_{\alpha}(k).$
Let $\mathop{}\!\mathrm{d}S_{\alpha,k}$ be the surface measure on $\sigma_{\alpha,k}$. Then
$\int_{M}\chi_{\alpha}\star|T_{\lambda}|=\int_{K_{\alpha}}\int_{\sigma_{\alpha,k}}\chi_{\alpha}\mathop{}\!\mathrm{d}S_{\alpha,k}\mathop{}\!\mathrm{d}\mu_{\alpha}(k),$
so summing in $\alpha$, we obtain
$\sum_{\alpha}\int_{K_{\alpha}}\int_{\sigma_{\alpha,k}}\chi_{\alpha}F\mathop{}\!\mathrm{d}\mu_{\alpha}(k)=\mathbf{M}(\lambda)=\sum_{\alpha}\int_{K_{\alpha}}\int_{\sigma_{\alpha,k}}\chi_{\alpha}\mathop{}\!\mathrm{d}S_{\alpha,k}\mathop{}\!\mathrm{d}\mu_{\alpha}(k).$ (3.3)
We claim that $\lambda$ is almost calibrated in the sense that for every $\alpha$ and $\mu_{\alpha}$-almost every $k$, $\sigma_{\alpha,k}$ is calibrated. If this is not true, then we may select $\beta$ and $K\subseteq K_{\beta}$ with $\mu_{\beta}(K)>0$, such that for every $k\in K$, $\int_{\sigma_{\beta,k}}F<\mathrm{vol}(\sigma_{\beta,k})$. Since $0<\chi_{\beta}\leq 1$ and $F/\mathop{}\!\mathrm{d}S_{\beta,k}\leq 1$ on $\sigma_{\beta,k}$, this is only possible if
$\int_{\sigma_{\beta,k}}\chi_{\beta}F<\int_{\sigma_{\beta,k}}\chi_{\beta}\mathop{}\!\mathrm{d}S_{\beta,k}.$
Integrating over $K$, and using the fact that in general we have $\int_{\sigma_{\alpha,k}}\chi_{\alpha}F\leq\int_{\sigma_{\alpha,k}}\chi_{\alpha}\mathop{}\!\mathrm{d}S_{\alpha,k}$, we conclude that
$\sum_{\alpha}\int_{K_{\alpha}}\int_{\sigma_{\alpha,k}}\chi_{\alpha}F\mathop{}\!\mathrm{d}\mu_{\alpha}(k)<\sum_{\alpha}\int_{K_{\alpha}}\int_{\sigma_{\alpha,k}}\chi_{\alpha}\mathop{}\!\mathrm{d}S_{\alpha,k}\mathop{}\!\mathrm{d}\mu_{\alpha}(k)$
which contradicts (3.3).
To upgrade $\lambda$ from an almost calibrated lamination to a calibrated lamination, we first, given $\sigma_{\alpha,k}$, choose $k_{j}$ such that $\sigma_{\alpha,k_{j}}$ is calibrated and $k_{j}\to k$. By Theorem 2.1, we can find a continuous $d-2$-form $A$ defined near $\sigma_{\alpha,k}$ with $F=\mathop{}\!\mathrm{d}A$. This justifies the following application of Stokes’ theorem:
$\int_{\sigma_{\alpha,k}}F=\int_{\partial\sigma_{\alpha,k}}A.$
Since $k_{j}\to k$, and $A$ is continuous,
$\mathrm{vol}(\sigma_{\alpha,k})$ $=\lim_{j\to\infty}\mathrm{vol}(\sigma_{\alpha,k_{j}})=\lim_{j\to\infty}\int_{\sigma_{\alpha,k_{j}}}F=\lim_{j\to\infty}\int_{\partial\sigma_{\alpha,k_{j}}}A=\int_{\partial\sigma_{\alpha,k}}A=\int_{\sigma_{\alpha,k}}F.$
To establish the converse, suppose that $\lambda$ is $F$-calibrated, and let notation be as above. Since $\lambda$ is $F$-calibrated, for every $\alpha$ and every $k$, the area form on $\sigma_{\alpha,k}$ is $F$. Therefore
$\int_{M}T_{\lambda}\wedge F$ $=\sum_{\alpha}\int_{K_{\alpha}}\int_{\sigma_{\alpha,k}}\chi_{\alpha}F\mathop{}\!\mathrm{d}\mu_{\alpha}(k)=\mathbf{M}(T_{\lambda}).\qed$
Lemma 3.4.
Suppose that $M$ is closed. Let $F$ be a calibration, and let $\lambda$ be a measured oriented $F$-calibrated lamination. Then:
  1. (1)
    $\lambda$ is homologically minimizing.
  2. (2)
    If $G$ is a calibration and cohomologous to $F$, then $\lambda$ is $G$-calibrated.
Proof.
Every leaf of $\lambda$ is $F$-calibrated, hence minimal. Since $\lambda$ is $F$-calibrated, so is $T_{\lambda}$ by Lemma 3.3, but then by (3.2), it follows that $T_{\lambda}$ is $G$-calibrated, and hence $\lambda$ is $G$-calibrated. Moreover, since $T_{\lambda}$ is $F$-calibrated, a calibration argument shows that $\lambda$ is homologically minimizing. ∎
3.2. Functions of least gradient
The natural “dual objects” to calibrations are functions of least gradient, which we now define.
We begin with some topological preliminaries. Let $M$ be a closed oriented Riemannian manifold of dimension $d$, and let $\tilde{M}\to M$ be the universal covering map. Any homomorphism
$\alpha:\pi_{1}(M)\to\mathbf{R}$
induces a homomorphism $\alpha:H_{1}(M,\mathbf{R})\to\mathbf{R}$. Thus $\alpha$ is an element of $H^{1}(M,\mathbf{R})$ and by Poincaré duality, we view it as an element of $H_{d-1}(M,\mathbf{R})$. Concretely, the following are equivalent for a function $u\in BV_{\mathrm{loc}}(\tilde{M},\mathbf{R})$:
  1. (1)
    $u$ is $\alpha$-equivariant, meaning that for every deck transformation $c\in\pi_{1}(M)$, and every $x\in\tilde{M}$,
    $u(cx)=u(x)+\alpha(c).$
  2. (2)
    $\mathop{}\!\mathrm{d}u$ descends to a $1$-current on $M$ whose cohomology class is the Poincaré dual of $\alpha$.
In either case we write $[\mathop{}\!\mathrm{d}u]=\alpha$, and write $\int_{M}\star|\mathop{}\!\mathrm{d}u|$ or $\mathbf{M}(\mathop{}\!\mathrm{d}u)$ to refer to the mass of the $1$-current that $\mathop{}\!\mathrm{d}u$ induces on $M$. If $[\mathop{}\!\mathrm{d}u]=0$ then we identify $u$ with the function that it induces on $M$.
Definition 3.5.
Let $u\in BV(\tilde{M},\mathbf{R})$ be a $\pi_{1}(M)$-equivariant function. Suppose that, for every $v\in BV(M,\mathbf{R})$,
$\int_{M}\star|\mathop{}\!\mathrm{d}u|\leq\int_{M}\star|\mathop{}\!\mathrm{d}u+\mathop{}\!\mathrm{d}v|.$
Then $u$ has least gradient.
An $\alpha$-equivariant function $u$ has least gradient iff $\int_{M}\star|\mathop{}\!\mathrm{d}u|$ is the stable norm, $\|\alpha\|_{1}$, of $\alpha$, which we defined in §2.3.
Lemma 3.6.
For each $\alpha\in H_{d-1}(M,\mathbf{R})$, there exists an $\alpha$-equivariant function of least gradient on $\tilde{M}$.
Proof.
The argument here is a standard application of the direct method of the calculus of variations, so we just sketch the proof. Let $(u_{n})$ be a sequence of $\alpha$-equivariant functions such that
$\lim_{n\to\infty}\int_{M}\star|\mathop{}\!\mathrm{d}u_{n}|=\|\alpha\|_{1}.$
This sequence is bounded in $BV_{\mathrm{loc}}(\tilde{M},\mathbf{R})$, so by Alaoglu’s theorem, it has a subsequence which converges in the weakstar topology of $BV_{\mathrm{loc}}$ to some function $u$ such that $\mathbf{M}(\mathop{}\!\mathrm{d}u)\leq\|\alpha\|_{1}$. By testing $\mathop{}\!\mathrm{d}u_{n}$ against smooth $d-1$-forms on $M$, we see that $[\mathop{}\!\mathrm{d}u]=\alpha$ and so $u$ has least gradient. ∎
Theorem 3.7.
Assume that $d\leq 7$. Let $u\in BV(\tilde{M},\mathbf{R})$ be a $\pi_{1}(M)$-equivariant function which is nonconstant. The following are equivalent:
  1. (1)
    $u$ has least gradient.
  2. (2)
    There is a homologically minimizing lamination $\lambda_{u}$ on $M$ such that:
    1. (a)
      $T_{\lambda_{u}}=\mathop{}\!\mathrm{d}u$.
    2. (b)
      Every leaf of $\lambda_{u}$ is a minimal hypersurface.
    3. (c)
      Every leaf of $\lambda_{u}$ pulls back to a union of subsets of $\tilde{M}$ of the form $\partial\{u>y\}$ or $\partial\{u<y\}$ for some $y\in\mathbf{R}$.
Proof.
If $u$ has least gradient, then [BackusCML, Theorem B] implies that there is a measured oriented lamination $\tilde{\lambda}_{u}$ of minimal hypersurfaces on $\tilde{M}$ whose leaves are level sets of $u$, and whose Ruelle-Sullivan current is $\mathop{}\!\mathrm{d}u$. Since $u$ is equivariant, $\tilde{\lambda}_{u}$ descends to a lamination $\lambda_{u}$ on $M$ such that $\mathbf{M}(\lambda_{u})=\mathbf{M}(\mathop{}\!\mathrm{d}u)$. Since $u$ has least gradient, $\lambda_{u}$ is homologically minimizing.
Conversely, if such a lamination exists, [BackusCML, Theorem B] implies that $u$ locally has least gradient and $\mathbf{M}(\mathop{}\!\mathrm{d}u)=\mathbf{M}(\lambda_{u})$, so $u$ has least gradient. ∎
Combining the above two results, we see that if $d\leq 7$, every nonzero class in $H_{d-1}(M,\mathbf{R})$ contains a homologically minimizing lamination.
3.3. Duality of calibrations and laminations
Recall from §2.3 the definition of the costable norm, $\|\cdot\|_{\infty}$. If $F$ is a calibration in a cohomology class $\rho$, then either $\|\rho\|_{\infty}=1$ (because $F$ minimizes its $L^{\infty}$ norm in $\rho$, and $\|F\|_{L^{\infty}}=1$), or $F$ calibrates no currents whatsoever. Conversely, if $\|\rho\|_{\infty}=1$, then by Alaoglu’s theorem, there is a calibration in $\rho$.
It is natural to ask if there is a continuous calibration in $\rho$, as was assumed in [bangert_cui_2017, Freedman_2016]. If $d=2$ one might try to generalize the argument of [Evans08] to obtain a Hölder continuous calibration, but if $d\geq 8$ then continuous calibrations need not exist [liu2023homologically]. The situation that $3\leq d\leq 7$ remains unclear. If $\mathrm{Ric}_{M}\geq 0$, then the Bochner argument shows that the harmonic representative of $\rho$ is a calibration; however, the Bochner argument actually shows that $M=\mathbf{S}^{1}\times N$ where $N$ is the calibrated hypersurface, so this is not very interesting.
In the setting of the Dirichlet problem for a domain on euclidean space, Mazón, Rossi, and Segura de León [Mazon14] proved that a $BV$ function has least gradient iff it is calibrated by some calibration. In fact, the same duality holds here, but in the equivariant setting the proof is trivial.
Lemma 3.8.
Let $u\in BV_{\mathrm{loc}}(\tilde{M},\mathbf{R})$ be an equivariant function. The following are equivalent:
  1. (1)
    $u$ has least gradient.
  2. (2)
    There exists a calibration $F$ on $M$ such that $\mathop{}\!\mathrm{d}u$ is $F$-calibrated.
Proof.
If $\mathop{}\!\mathrm{d}u$ is $F$-calibrated, then we have by Stokes’ theorem and (2.2) that for any $v\in BV(M,\mathbf{R})$,
$\int_{M}\star|\mathop{}\!\mathrm{d}u|=\int_{M}\mathop{}\!\mathrm{d}u\wedge F=\int_{M}(\mathop{}\!\mathrm{d}u+\mathop{}\!\mathrm{d}v)\wedge F\leq\int_{M}\star|\mathop{}\!\mathrm{d}u+\mathop{}\!\mathrm{d}v|,$
so $u$ has least gradient.
Conversely, if $u$ has least gradient, then let $\alpha:=[\mathop{}\!\mathrm{d}u]$ and choose $\rho\in H^{d-1}(M,\mathbf{R})$ such that $\langle\rho,\alpha\rangle=\|\alpha\|_{1}$ and $\|\rho\|_{\infty}=1$. In particular, there exists a calibration $F$ such that $[F]=\rho$, and
$\int_{M}\mathop{}\!\mathrm{d}u\wedge F=\langle\rho,\alpha\rangle=\|\alpha\|_{1}=\int_{M}\star|\mathop{}\!\mathrm{d}u|,$
so that $u$ has least gradient. ∎
The above proof motivates the introduction of the following terminology from convex geometry. A flat in the stable unit sphere $\partial B$ is the intersection of $\partial B$ with a hyperplane. In particular, every flat is convex. If $\|\rho\|_{\infty}=1$, its dual flat is
$\rho^{*}:=\{\alpha\in\partial B:\langle\rho,\alpha\rangle=1\}.$
This set is convex, compact, and nonempty; in general it does not have to be a singleton. Every hyperplane in $H_{d-1}(M,\mathbf{R})$ takes the form $\{\alpha\in H_{d-1}(M,\mathbf{R}):\langle\rho,\alpha\rangle=t\}$ for some $\rho$ in the costable unit sphere and some $t\in\mathbf{R}$, so every flat in $\partial B$ is contained in $\rho^{*}$ for some $\rho\in\partial B^{*}$.
The next lemma was observed by Bangert and Cui, [bangert_cui_2017], in the setting that $F$ is continuous and we require no regularity on the laminations involved.
Lemma 3.9.
Suppose that $M$ is a closed Riemannian manifold of dimension $d\leq 7$. Let $\rho\in H^{d-1}(M,\mathbf{R})$ satisfy $\|\rho\|_{\infty}=1$, and let $F$ be a calibration in $\rho$. Then there exists an $F$-calibrated measured oriented lamination.
Proof.
Choose $\alpha\in\rho^{*}$, and let $u$ be an $\alpha$-equivariant function of least gradient. Then $\mathop{}\!\mathrm{d}u$ is $F$-calibrated, so the measured oriented lamination $\kappa$ given by Theorem 3.7 is $F$-calibrated by Lemma 3.3. ∎
In view of the Mazón–Rossi–Segura de León theorem and Lemma 3.9, it is natural to conjecture that if $F$ minimizes its $L^{\infty}$ norm subject to a boundary condition on a domain $U$ in euclidean space and $\|F\|_{L^{\infty}}=1$, then $F$ calibrates some function on $U$. The following example shows that this conjecture is false.
Example 3.10.
Let
$v(x+iy):=\arctan\left(\frac{y}{x}\right)$
defined on the open disk $U$ bounded by the circle $(x-2)^{2}+y^{2}=1$. Then $v$ is $\infty$-harmonic, meaning that
$\langle\nabla^{2}v,\nabla v\otimes\nabla v\rangle=0.$
To see this, it is best to work in polar coordinates, $x+iy=re^{i\theta}$. Then $v(re^{i\theta})=\theta$, so $\mathop{}\!\mathrm{d}v=\mathop{}\!\mathrm{d}\theta$. The euclidean metric is
$g=\mathop{}\!\mathrm{d}r^{2}+r^{2}\mathop{}\!\mathrm{d}\theta^{2},$
so the Christoffel symbol ${\Gamma^{\theta}}_{\theta\theta}$ vanishes. Then we compute
$\langle\nabla\mathop{}\!\mathrm{d}\theta,\mathop{}\!\mathrm{d}\theta\otimes\mathop{}\!\mathrm{d}\theta\rangle=\langle\nabla\mathop{}\!\mathrm{d}\theta,\partial_{\theta}\otimes\partial_{\theta}\rangle r^{-4}=r^{-4}{\Gamma^{\theta}}_{\theta\theta}=0.$
Also, $|\mathop{}\!\mathrm{d}\theta|=r^{-1}$, which only attains its maximum at the boundary point $x+iy=1$. In particular, $\|\mathop{}\!\mathrm{d}v\|_{L^{\infty}}=1$ and $\mathop{}\!\mathrm{d}v$ is a calibration on $U$. Since $v$ is $\infty$-harmonic, $v$ minimizes its Lipschitz constant, $\|\mathop{}\!\mathrm{d}v\|_{L^{\infty}}$, among all functions with the same boundary data [Crandall2008]. But if $u$ is a function on $U$ such that $\mathop{}\!\mathrm{d}u$ is $\mathop{}\!\mathrm{d}v$-calibrated, then
$\operatorname{supp}\mathop{}\!\mathrm{d}u\subseteq\{|\mathop{}\!\mathrm{d}v|=1\}\subset\partial U$
so $u$ is constant away from the boundary, hence is constant.
A more geometric way to visualize this phenomenon is to notice that the streamlines of $v$ – that is, the integral curves of the gradient of $v$ – are the circles centered on 00. If $u$ was $\mathop{}\!\mathrm{d}v$-calibrated, then the level sets of $u$ would correspond to the streamlines of $v$. However, since any $\mathop{}\!\mathrm{d}v$-calibrated function $u$ has least gradient, the level sets of $u$ must be straight lines.
4. Construction of the canonical lamination
Throughout this section, we fix a closed oriented Riemannian manifold $M$ of dimension $2\leq d\leq 7$, and a cohomology class $\rho\in H^{d-1}(M,\mathbf{R})$ in the costable unit sphere: $\|\rho\|_{\infty}=1$. We prove Theorem 1.2: the set of complete immersed hypersurfaces, which are calibrated by every calibration in $\rho$, is the set of leaves of a lamination with Lipschitz regularity.
Let $F$ be a calibration in $\rho$. The set $S:=\{|F|=1\}$ need not be the support of a lamination $\lambda$; and even if it was, we would not be able to conclude that $F$ calibrates $\lambda$. For example, if $d\geq 3$, then one can exploit the possible nonintegrability of $\star F$ to produce counterexamples [bangert_cui_2017, §4]. More starkly, if $d=2$, then the main theorem of [daskalopoulos2020transverse] then implies that $S$ contains a geodesic lamination $\lambda$; on the other hand, the main theorem of [backus2024lipschitz] implies that any closed set containing $\operatorname{supp}\lambda$ can be realized as the set $\{|G|=1\}$ for some calibration $G$ in $\rho$. If $M$ is hyperbolic, then $\lambda$ has Hausdorff dimension $1$ [Birman1985], so “almost every” closed subset of $M$ is $\{|G|=1\}$ for some $G$.
We shall construct a lamination $\lambda_{F}$ whose support is contained in $S$, such that every $F$-calibrated hypersurface is a leaf of $\lambda_{F}$. By Theorem 2.8, we must establish the following:
  1. (1)
    There is an $F$-calibrated hypersurface.
  2. (2)
    There is a uniform bound on the curvatures of the $F$-calibrated hypersurfaces.
  3. (3)
    Any two $F$-calibrated hypersurfaces are disjoint.
  4. (4)
    The limit of a sequence of $F$-calibrated hypersurfaces is a $F$-calibrated hypersurface.
The nontriviality condition (1) is nothing more than Lemma 3.9. The hypersurface furnished by that lemma is actually calibrated by every calibration in $\rho$, so the intersection of all laminations $\lambda_{F}$ is nonempty; this intersection shall be the canonical calibrated lamination.
We now show that the leaves of the putative canonical lamination satisfy the necessary curvature bounds; this is a little subtle because the leaves are injectively immersed but not embedded. In the below lemmata, let $r_{*}$ be the minimum of the injectivity radius of $M$ and $\delta\|\mathrm{Riem}_{M}\|_{C^{0}}^{-1/2}$, where $\delta>0$ is a dimensional constant to be determined later. Let $\mathbf{S}^{d-1}$ be the round sphere of dimension $d-1$.
Lemma 4.1.
Let
$U:=\bigcup_{x\in M}\{\xi\in T_{x}M:0<|\xi|<r_{*}\},$
and let $F:U\to M$ be the exponential map. Then for every injectively immersed hypersurface $N\subset M$, $F$ is transverse to $N$.
Proof.
We must show that for every $(x,\xi)\in U$ such that $F(x,\xi)\in N$, the image of
$\mathop{}\!\mathrm{d}F(x,\xi):T_{(x,\xi)}T_{x}M\to T_{F(x,\xi)}M$
contains a vector not tangent to $N$. Let $\eta$ be the unit normal to $N$ at $F(x,\xi)$, and let $\overline{\eta}$ be the parallel transport of $\eta$ along the unique geodesic $\gamma$ from $F(x,\xi)$ to $x$. Viewing $\overline{\eta}$ as an element of $T_{F(x,\xi)}T_{x}M$, we see that if $\delta$ was chosen small enough, then $\mathop{}\!\mathrm{d}F(x,\xi)\overline{\eta}$ lies in a small neighborhood of $\eta$. Indeed, if $\delta$ was chosen small enough, then $\gamma$ is much shorter than the curvature scale $\|\mathrm{Riem}_{M}\|_{C^{0}}^{-1/2}$. In particular, $\mathop{}\!\mathrm{d}F(x,\xi)\overline{\eta}$ is not tangent to $N$. ∎
Lemma 4.2.
For every calibration $F$, every complete injectively immersed $F$-calibrated hypersurface $N\subset M$, every $x\in M$, every $0<r\leq r_{*}$, and every component $N^{\prime}$ of $N\cap B(x,r)$,
$\mathrm{vol}(\mathbf{B}^{d-1})\leq\mathrm{vol}(N^{\prime})\leq 2\mathrm{vol}(\mathbf{S}^{d-1})r^{d-1}.$ (4.1)
Proof.
If $\delta$ was chosen small enough, then
$\mathrm{vol}(\partial B(x,r))<2\mathrm{vol}(\mathbf{S}^{d-1})r^{d-1}.$
Let $F:U\to M$ be the exponential map as in Lemma 4.1, so that $F$ is transverse to $N$. By putting polar coordinates on each tangent space, we may view $U$ as a fiber bundle,
$M\times(0,r_{*})\to U\to\mathbf{S}^{d-1}.$
By the Thom transversality theorem, for almost every $(x,r)\in M\times(0,r_{*})$, the induced map
$f_{x,r}:\mathbf{S}^{d-1}$ $\to M$
$\omega$ $\mapsto F(x,r\omega)$
is transverse to $N$. But $f_{x,r}$ is the embedding $\mathbf{S}^{d-1}\to\partial B(x,r)$. The estimate (4.1) is preserved by slight perturbations of $r$, so we may use the above considerations to reduce to the case that $N$ is transverse to $\partial B(x,r)$.
Let $N^{\prime}$ be a component of $N\cap B(x,r)$, so that $N^{\prime}$ is embedded (not just injectively immersed). By transversality, $N^{\prime}\cap\partial B(x,r)$ is diffeomorphic to a closed $d-2$-dimensional submanifold of $\mathbf{S}^{d-1}$. Since $H_{d-2}(\mathbf{S}^{d-1},\mathbf{R})=0$, there exists a relatively open set $V\subseteq\partial B(x,r)$ which is bounded by $N\cap\partial B(x,r)$. Because of how we chose $r_{*}$, we may use the $L^{\infty}$ Poincaré lemma, Theorem 2.1, to find a continuous $d-2$-form $A$ on a neighborhood of the closure of $B(x,r)$, such that $F=\mathop{}\!\mathrm{d}A$. Then
$\mathrm{vol}(N\cap B(x,r))$ $=\int_{N\cap B(x,r)}F=\int_{N\cap\partial B(x,r)}A=\int_{V}F\leq\mathrm{vol}(V)\leq\mathrm{vol}(\partial B(x,r))$
$<2\mathrm{vol}(\mathbf{S}^{d-1})r^{d-1}.\qed$
Lemma 4.3.
There exists a constant $C>0$, only depending on $M$, such that for every calibration $F$ and complete injectively immersed $F$-calibrated hypersurface $N$, we have the curvature bound
$\|\mathrm{I\!I}_{N}\|_{C^{0}}\leq C.$ (4.2)
Proof.
Let $x\in N$ and let $r>0$ be small enough depending on $M$. Then each component $N^{\prime}$ of $N\cap B(x,r)$ is $F$-calibrated, and therefore a stable minimal hypersurface. By (4.1), $\mathrm{vol}(N^{\prime})\lesssim r^{d-1}$. So by [Schoen81, pg785, Corollary 1] (see also [colding2011course, Chapter 2, §§4-5]),
$\|\mathrm{I\!I}_{N^{\prime}}\|_{C^{0}(B(x,r/2))}\lesssim_{d,\|\mathrm{Riem}_{g}\|_{C^{0}(B(x,2r))}}\frac{1}{r}.$
Since $N^{\prime}$ was an arbitrary component, the same estimate holds for $N$. Using the compactness of $M$, we may cover it by finitely many balls in which estimates of this form hold to conclude (4.2). ∎
Lemma 4.4.
Let $F$ be a calibration, and let $N,N^{\prime}$ be immersed $F$-calibrated hypersurfaces. Then:
  1. (1)
    If $N\cap N^{\prime}$ is nonempty, then for each $x\in N\cap N^{\prime}$ there is an open neighborhood $U$ of $x$ such that $N\cap U=N^{\prime}\cap U$.
  2. (2)
    If $N\cap N^{\prime}$ is nonempty, and $N,N^{\prime}$ are complete and connected, then $N=N^{\prime}$.
  3. (3)
    $N$ is injectively immersed.
Proof.
We first observe that for each $x\in N$, $(\star F(x))^{\sharp}$ is the (unique) normal vector to $N$ at $x$ (and similarly for $N^{\prime}$), and so if $x\in N\cap N^{\prime}$ then $N,N^{\prime}$ have the same tangent space at $x$. So for each $x\in N\cap N^{\prime}$, there exists $r>0$ and normal coordinates $(\xi,\eta)\in\mathbf{R}^{d-1}\times\mathbf{R}$ on $B(x,r)$ based at $x$, such that for each pair of sheets $N_{*}\subseteq N\cap B(x,r)$, $N_{*}^{\prime}\subseteq N^{\prime}\cap B(x,r)$ which contain $x$, there exists a relatively open set $V\subseteq\{\eta=0\}$, an open set $U\subseteq B(x,r)$ containing $x$, and functions $u,u^{\prime}:V\to\mathbf{R}$ such that:
  1. (0–1)
    $N_{*}\cap U=\{(\xi,u(\xi)):\xi\in V\}$.
  2. (0–2)
    $N_{*}^{\prime}\cap U=\{(\xi,u^{\prime}(\xi)):\xi\in V\}$.
  3. (0–3)
    $u(0)=u^{\prime}(0)=0$.
  4. (0–4)
    If $u(\xi)=u^{\prime}(\xi)$ then $\mathop{}\!\mathrm{d}u(\xi)=\mathop{}\!\mathrm{d}u^{\prime}(\xi)$.
Let $v:=u-u^{\prime}$. Then:
  1. (0–1)
    $v(0)=0$.
  2. (0–2)
    $v$ satisfies a linear elliptic PDE on $V$ [colding2011course, Proof of Theorem 7.3].
  3. (0–3)
    If $v(\xi)=0$ then $\mathop{}\!\mathrm{d}v(\xi)=0$.
We claim that $v$ is identically 00. If this is not true, the set $\{v=0\}=\{v=\mathop{}\!\mathrm{d}v=0\}$ is $d-3$-rectifiable [Hardt89, Lemma 1.9], but $\dim V=d-1$, so $\{v\neq 0\}$ is connected. So either $v\geq 0$ or $v\leq 0$, and $v$ has a zero; this contradicts the maximum principle.
The above discussion shows that $N_{*}\cap U=N_{*}^{\prime}\cap U$. Taking $N=N^{\prime}$ we see that $N$ only has one sheet in $B(x,r)$ which contains $x$, so (3) holds. So running the same argument, without assuming that $N=N^{\prime}$, yields (1). A continuity argument then implies (2). ∎
We must show that a limit of $F$-calibrated hypersurfaces is $F$-calibrated, and to make this precise we shall need the notion of a Vietoris limit superior of a sequence of closed sets, [kechris2012classical, §4.F]. If $(K_{n})$ is a sequence of closed subsets of $M$, then $\limsup_{n\to\infty}K_{n}$ is the set of all $x$ such that for every open set $U\ni x$, there exist infinitely many $n\in\mathbf{N}$ such that $U\cap K_{n}\neq\emptyset$; one easily checks that $\limsup_{n\to\infty}K_{n}$ is closed.
Lemma 4.5.
Let $F$ be a calibration, let $(N_{n})$ be a sequence of $F$-calibrated complete connected immersed hypersurfaces, and let $K:=\limsup_{n\to\infty}\overline{N_{n}}$. For every $x\in K$ there exists a $F$-calibrated complete connected immersed hypersurface $N\subseteq K$ such that $x\in N$.
Proof.
By taking a subsequence, we may assume that there exist $x_{n}\in N_{n}$ such that $x_{n}\to x$. By Lemma 4.4, we may also assume that if $N_{n}\cap N_{m}$ is nonempty then $n=m$. Combining this with the curvature bound (4.2), we obtain the hypotheses of [BackusCML, Lemma 2.4]. The conclusion of that lemma is that for every $\delta>0$ there exists $r>0$ only depending on $M$, and normal coordinates $(\xi,\eta)\in\mathbf{R}^{d-1}\times\mathbf{R}$ on $B(x,r)$ based at $x$ such that for every $n$,
$\|\mathbf{n}_{N_{n}}-\partial_{\eta}\|_{C^{0}(B(x,r))}\leq\delta.$ (4.3)
If $\delta$ was chosen small enough, depending only on $M$, then by the vertical line test, there exists a relatively open set $V\subseteq\{\eta=0\}$ and a sequence of functions $u_{n}$ on $V$, such that:
  1. (0–1)
    $x\in V$.
  2. (0–2)
    There exists $c_{0}>0$ which only depends on $M$ such that $\mathrm{diam}(V)\geq c_{0}$.
  3. (0–3)
    For every $n$, $N_{n}\cap\{(\xi,\eta)\in B(x,r):\xi\in V\}=\{(\xi,u_{n}(\xi)):\xi\in V\}$.
The functions $u_{n}$ solve the minimal surface equation,
$Pu(\xi):=F(\xi,u(\xi),\mathop{}\!\mathrm{d}u(\xi),\nabla^{2}u(\xi))=0$
where one can use [colding2011course, (7.21)] to show that $F$ has the form
$F(\xi,\eta,A,B):=\operatorname{tr}B+O((|\xi|+|\eta|+|A|)(1+|B|))$
where the implied constant only depends on $M$. But $|\xi|+|u_{n}(\xi)|\lesssim r$ and, if
$\|\mathbf{n}_{N_{n}}-\partial_{\eta}\|_{C^{0}(B(x,r))}\leq\frac{1}{10},$
then one may show that
$|\mathop{}\!\mathrm{d}u_{n}(\xi)|\leq\|\mathop{}\!\mathrm{d}u_{n}\|_{C^{0}}\lesssim\|\mathbf{n}_{N_{n}}-\partial_{\eta}\|_{C^{0}(B(x,r))}.$
So by (4.3), we can first choose $\delta$ small enough depending on $M$, and then choose $r$ small enough depending on $\delta$, so that for every $n$ large enough depending on $r$, the equation $Pu_{n}=0$ is uniformly elliptic. In particular, by the interior Schauder estimate [gilbarg2015elliptic, Theorem 6.2], we may choose $\delta$ small enough, depending only on $M$, that there exists a connected, relatively open set $W\subseteq V$ such that:
  1. (0–1)
    $x\in W$.
  2. (0–2)
    There exists $c_{1}>0$ which only depends on $M$ such that $\mathrm{diam}(W)\geq c_{1}$.
  3. (0–3)
    For every sufficiently large $n$, $\|u_{n}\|_{C^{3}(W)}\leq 1$.
Therefore there exists $u\in C^{2}(W)$ such that:
  1. (0–1)
    After passing to a subsequence, $u_{n}\to u$ in $C^{2}(W)$.
  2. (0–2)
    $N^{x}:=\{(\xi,u(\xi)):\xi\in W\}$ contains $x$.
  3. (0–3)
    $N^{x}\subseteq K$.
We moreover claim that, possibly after shrinking $W$ (while preserving (0–1) and (0–2)):
  1. (0–1)
    $N^{x}$ is $F$-calibrated.
  2. (0–2)
    $N^{x}$ is geodesically convex.
  3. (0–3)
    There exists $c_{2}>0$ which only depends on $M$ such that $\operatorname{dist}_{N^{x}}(x,\partial N^{x})\geq c_{2}$.
To prove this, let $N_{n}^{x}:=\{(\xi,u_{n}(x)):\xi\in W\}$. If $\mathrm{diam}(W)$ was chosen small enough (depending only on $M$), then we can use the $L^{\infty}$ Poincaré lemma, Theorem 2.1, to find a continuous $d-2$-form $A$ on $W$ such that $\mathop{}\!\mathrm{d}A=F$. Since $u_{n}\to u$ in $C^{2}(W)$, we can compute using Stokes’ theorem
$\int_{N^{x}}F=\int_{\partial N^{x}}A=\lim_{n\to\infty}\int_{\partial N_{n}^{x}}A=\lim_{n\to\infty}\int_{N_{n}^{x}}F=\lim_{n\to\infty}\mathrm{vol}(N_{n}^{x})=\mathrm{vol}(N^{x}),$
which proves (0–1). By shrinking $W$ slightly more, we can impose (0–2). Moreover, since $\partial N^{x}\subset\partial W$, and the curvature bound (4.2) allows us to compare distances in $M$ and distances in $N^{x}$, (0–3) follows from (0–2).
Let $N$ be the union of all $F$-calibrated connected immersed hypersurfaces contained in $K$ which extend $N^{x}$. If $N$ is incomplete, then there exists $y\in N$ such that $\operatorname{dist}_{N}(y,\partial N)<c_{2}$. Then $y\in K$, so there exists a $F$-calibrated connected immersed hypersurface $N^{y}\subseteq K$ such that $y\in N^{y}$ and (0–2) and (0–3) hold. But $N^{y}\subseteq N$, so by (0–2) and (0–3),
$\operatorname{dist}_{N}(y,\partial N^{y})\geq\operatorname{dist}_{N}(y,\partial N^{y})\geq c_{2},$
which is a contradiction. Therefore $N$ is complete. ∎
Lemma 4.6.
Let $F$ be a calibration in $\rho$. Then the set of $F$-calibrated connected complete immersed hypersurfaces is the set of leaves of a lamination $\lambda_{F}$, which contains every measured oriented $F$-calibrated lamination.
Proof.
Let $\mathscr{L}_{F}$ be the set of connected complete immersed $F$-calibrated hypersurfaces. By Lemma 4.4, $\mathscr{L}_{F}$ consists of pairwise disjoint injectively immersed minimal hypersurfaces. The curvature bound (4.2) only depends on $M$, and implies that the elements of $\mathscr{L}_{F}$ have curvatures bounded uniformly in $C^{0}$. By Lemma 3.9, $\mathscr{L}_{F}$ is nonempty.
Let $E$ be the union of all elements of $\mathscr{L}_{F}$. If $(x_{n})$ is a sequence in $E$, say $x_{n}\in N_{n}$ for some $N_{n}\in\mathscr{L}_{F}$, and $x_{n}\to x$, then $x\in\limsup_{n\to\infty}\overline{N_{n}}$. So by Lemma 4.5, there exists $N\in\mathscr{L}_{F}$ such that $x\in N$. In particular, $x\in E$, so $E$ is closed.
By the above discussion and Theorem 2.8, $\mathscr{L}_{F}$ is the set of leaves of some lamination $\lambda_{F}$. ∎
Proof of Theorem 1.2.
Let $S$ be the set of calibrations in $\rho$, which is nonempty since $\|\rho\|_{\infty}=1$. Then there is a lamination which is $F$-calibrated by every $F\in S$. Indeed, by Lemma 3.9, there is a measured oriented lamination $\kappa$ which is $F$-calibrated for some $F\in S$, and by Lemma 3.4, $\kappa$ is $F$-calibrated for every $F\in S$.
For every $F\in S$, let $\lambda_{F}$ be the calibrated lamination produced by Lemma 4.6. By Lemma 2.11, there is a lamination $\lambda_{\rho}$ whose set of leaves is the intersection over $F\in S$ of the sets of leaves of $\lambda_{F}$. Then $\lambda_{\rho}$ has all desired properties. ∎
5. Transverse measures on the canonical lamination
5.1. Ergodic theory of $\lambda_{\rho}$
Let $M$ be a closed oriented Riemannian manifold of dimension $2\leq d\leq 7$. For each oriented lamination $\lambda$ in $M$, let $\mathcal{M}(\lambda)$ be the set of transverse probability measures to sublaminations of $\lambda$. This set inherits the vague topology on the space of Borel probability measures on $\operatorname{supp}\lambda$; in view of (2.3), this topology is the same as the topology on the space of measured laminations (see [BackusCML]) restricted to $\mathcal{M}(\lambda)$. It is clear that $\mathcal{M}(\lambda)$ is convex, and one may use the compactness of the space of Borel probability measures on the compact metrizable space $\operatorname{supp}\lambda$ [kechris2012classical, Theorem 17.23] to show that $\mathcal{M}(\lambda)$ is compact. By the Krein-Milman theorem, if $\mathcal{M}(\lambda)$ is nonempty, then so is its set of extreme points, $\mathcal{E}(\mathcal{M}(\lambda))$.
Definition 5.1.
A measure $\mu\in\mathcal{M}(\lambda)$ is ergodic if, for every saturated set $E$, either $\mu(E)=0$ or $\mu(E)=1$.
Lemma 5.2.
Every extreme point of $\mathcal{M}(\lambda)$ is ergodic, and the set of ergodic measures is linearly independent in the space of signed Borel measures on $\operatorname{supp}\lambda$.
Proof.
The first claim is an easy modification of the proof of [einsiedler2010ergodic, Theorem 4.4], and the second is essentially the proof that every ultrafilter on a finite set is principal. To be more precise, let $S$ be a finite set of ergodic measures, and choose $c_{\mu}\in\mathbf{R}$ such that $\sum_{\mu\in S}c_{\mu}\mu=0$. The measures in $S$ are determined by their values on saturated sets, so if some coefficient $c_{\nu}$ is nonzero, then there exists a saturated set $E$ and a proper subset $T\subset S$ such that:
  1. (0–1)
    $\nu\in T$.
  2. (0–2)
    For every $\mu\in T$, $\mu(E)=1$.
  3. (0–3)
    For every $\mu\in S\setminus T$, $\mu(E)=0$.
Then $S^{\prime}:=\{1_{E}\mu:\mu\in T\}$ satisfies:
  1. (0–1)
    For every $\mu\in S^{\prime}$, $\mu$ is ergodic.
  2. (0–2)
    $\sum_{\mu\in S^{\prime}}c_{\mu}\mu=0$.
  3. (0–3)
    There exists $\nu\in S^{\prime}$ such that $c_{\nu}$ is nonzero.
  4. (0–4)
    $\operatorname{card}S^{\prime}<\operatorname{card}S$.
Therefore we can repeat the argument with $S$ replaced by $S^{\prime}$. After finitely many iterations, we reduce to the case that $\operatorname{card}S\leq 1$, in which case we have a contradiction. ∎
Now let $B$ be the stable unit ball of $H_{d-1}(M,\mathbf{R})$ and $B^{*}$ be the costable unit ball. For each $\rho\in B^{*}$, we consider the set $\mathcal{M}(\lambda_{\rho})$ of transverse probability measures to the canonical lamination, $\lambda_{\rho}$. By Lemma 3.4, $\mathcal{M}(\lambda_{\rho})$ is the set of measured oriented laminations which are calibrated by some calibration in $\rho$. The map which sends a measured oriented lamination to its homology class restricts to a an affine map $\mathcal{M}(\lambda_{\rho})\to\rho^{*}$, which is surjective by Lemma 3.9. In particular, if $\alpha$ is an extreme point of the dual flat, $\rho^{*}$, then $\alpha$ is the homology class of an ergodic measure.
We summarize the above discussion as Corollary 1.3.
5.2. The Arnoux–Levitt lemma
Following an idea of Auer and Bangert [Auer01], we study $\mathcal{M}(\lambda_{\rho})$ using an ergodic-theoretic lemma of Arnoux and Levitt, [Arnoux1986, Proposition 3.1]. We need a more general version of the Arnoux–Levitt lemma, and the original proof is in French, so we include a proof here.
Let us identify transverse measures with positive transverse cocycles (cocycles which act on curves transverse to the lamination and are cooriented with the lamination); this is standard, and we refer to [daskalopoulos2020transverse, §7.2] for a justification of this identification.
Lemma 5.3.
Let $\lambda$ be an oriented lamination, $U\subseteq M$ open, and $\mu,\nu\in\mathcal{M}(\lambda)$. Assume that:
  1. (1)
    $\lambda$ is not a closed hypersurface, and there is a leaf of $\lambda$ which is dense in $\operatorname{supp}\lambda\cap U$.
  2. (2)
    $\mu(U)=\nu(U)=1$.
  3. (3)
    There exists $b\geq 0$ such that for every $1$-cycle $C\subset U$ which is transverse to $\lambda$, $\mu(C)-\nu(C)\in b\mathbf{Z}$.
Then $\mu=\nu$.
Proof.
We follow [Arnoux1986, §3, Lemme] which is a similar result when $\lambda$ is a minimal component of a foliation.
Let $C\subset U$ be a transverse curve to $\lambda$, which is cooriented with $\lambda$, and such that $\mu(C)>0$ and $\nu(C)>0$. By assumptions (1) and (2), there exists a leaf $N$ such that:
  1. (0–1)
    $N$ is dense in $\operatorname{supp}\lambda\cap U$.
  2. (0–2)
    $N\cap C$ is infinite and dense in $\operatorname{supp}\lambda\cap U\cap C$.
If we prove that $\mu(C)=\nu(C)$ for every sufficiently short cooriented transverse curve $C$, then the result follows for all curves by $\sigma$-additivity. Therefore we may shorten $C$ so that:
  1. (0–1)
    $C$ begins and ends on $N$.
  2. (0–2)
    If $b>0$ then $\mu(C)+\nu(C)<b$.
Let $\sigma\subset N$ be a curve from the beginning of $C$ to the end of $C$, and let $C^{\prime}$ be a deformation of $C\cup\sigma$ through homotopies which leave $\lambda\setminus N$ fixed, so that $C^{\prime}$ is a transverse cycle to $\lambda$. Then, by (3), for some $k\in\mathbf{Z}$,
$\mu(C)=\mu(C^{\prime})=\nu(C^{\prime})+kb=\nu(C)+kb.$
If $b=0$ then we are done; otherwise, since $\mu(C)+\nu(C)<b$, it follows that $k=0$. ∎
Lemma 5.4 (Arnoux–Levitt lemma).
Let $\lambda$ be an oriented lamination, and let $\mathcal{I}(\lambda)$ be the set of ergodic probability measures transverse to sublaminations $\kappa\subseteq\lambda$ such that $\kappa$ is not a closed hypersurfaces. Then the homology classes of measures in $\mathcal{I}(\lambda)$ are linearly independent, and if they span $H_{d-1}(M,\mathbf{R})$, then $H_{d-1}(M,\mathbf{R})=0$.
Proof.
Let $\kappa_{1},\dots,\kappa_{q}$ be distinct sublaminations of $\lambda$, such that for each $1\leq i\leq q$:
  1. (0–1)
    There exists $m_{i}\geq 1$ and distinct probability measures $\mu_{i}^{1},\dots,\mu_{i}^{m_{i}}$ such that for each $1\leq j\leq m_{i}$, $\mu_{i}^{j}$ is ergodic and transverse to $\kappa_{i}$.
  2. (0–2)
    $\kappa_{i}$ is not a closed hypersurface.
Notice that if $q>0$ then $H_{d-1}(M,\mathbf{R})\neq 0$. We define open sets $U_{i}$ and leaves $N_{i}$ of $\kappa_{i}$, such that:
  1. (0–1)
    $N_{i}$ is dense in $\operatorname{supp}\kappa_{i}\cap U_{i}$.
  2. (0–2)
    For every $j$, $\mu_{i}^{j}(U_{i})=1$.
To do this we break into cases:
  1. (0–1)
    Suppose that $\kappa_{i}$ is not a foliation with a dense leaf. By the Morgan–Shelan decomposition (Theorem 2.12) and the ergodicity of $\mu_{1}^{i}$, $\kappa_{i}$ is exceptional, and if $i^{\prime}\neq i$ then $\operatorname{supp}\kappa_{i^{\prime}}$ avoids an open set $U_{i}$ containing $\operatorname{supp}\kappa_{i}$. Let $N_{i}$ be any leaf of $\kappa_{i}$.
  2. (0–2)
    Suppose that $\kappa_{i}$ is a foliation with a dense leaf $N_{i}$. In particular $\kappa_{i}=\lambda$. For any $i^{\prime}\neq i$, and any $j$, $\mu_{i}^{j}$ is ergodic and $\operatorname{supp}\kappa_{i^{\prime}}$ is saturated, so $\mu_{i}^{j}(\operatorname{supp}\kappa_{i^{\prime}})=0$. Therefore $U_{i}:=M\setminus\bigcup_{i^{\prime}\neq i}\operatorname{supp}\kappa_{i}$ has $\mu_{i}^{j}(U_{i})=1$. Since $\kappa_{i^{\prime}}$ is not a foliation, $\operatorname{supp}\kappa_{i^{\prime}}$ is a saturated closed subset of $M$ which is not $M$; therefore it misses the dense leaf $N_{i}$.
Next we show:
  1. (0–1)
    For every $i$, $([\mu_{i}^{j}])_{j}$ is linearly independent.
Suppose that there are $a_{j}\in\mathbf{R}$ such that $\sum_{j}a_{j}[\mu^{i}_{j}]=0$, let $\mu$ be the sum of $a_{j}\mu_{i}^{j}$ over $j$ such that $a_{j}>0$, and let $\nu$ be the sum of $-a_{j}\mu_{i}^{j}$ over $j$ such that $a_{j}<0$. Then $[\mu]-[\nu]=0$, so by Lemma 5.3 with $b=0$ and $U=M$, $\mu-\nu=0$, hence $\sum_{j}a_{j}\mu^{i}_{j}=0$. By ergodicity, $(\mu_{i}^{j})_{j}$ is linearly independent, so $a_{j}=0$, establishing (0–1).
We claim there are $1$-cocycles $t_{i}$ such that:
  1. (0–1)
    There exists a $1$-cycle $C_{i}\subset U_{i}$ such that $t_{i}(C_{i})\neq 0$.
  2. (0–2)
    For every $j\neq i$, and every $1$-cycle $C\subset U_{j}$, $t_{i}(C)=0$.
  3. (0–3)
    There exists $q_{i}\in\mathbf{Q}$ such that for every $1$-cycle $C$, $t_{i}(C)\in q_{i}\mathbf{Z}$.
By composing with the natural homomorphism $H_{1}(M,\mathbf{Z})\to H_{1}(M,\mathbf{R})$, we can think of the cohomology class of $\mu_{i}^{1}$ as a homomorphism
$[\mu_{i}^{1}]:H_{1}(M,\mathbf{Z})\to\mathbf{R}.$
Since $M$ is compact, $H_{1}(M,\mathbf{Z})$ is finitely generated and so we can slightly perturb the value of $[\mu_{i}^{1}]$ on the generators to obtain $t_{i}$ with the desired properties.
To complete the proof it is enough to show that
  1. (0–1)
    $(t_{i},[\mu_{i}^{j}])_{i,j}$ is linearly independent.
Suppose that there are $a_{i}^{j},a_{i}\in\mathbf{R}$ such that
$\sum_{i,j}a_{i}^{j}[\mu_{i}^{j}]+\sum_{i}a_{i}t_{i}=0.$
Then for every $1$-cycle $C$ in $U_{i}$,
$\sum_{j}a_{i}^{j}\mu_{i}^{j}(C)=-a_{i}t_{i}(C)\in a_{i}q_{i}\mathbf{Z}.$
Let $\mu_{i}^{+}$ be the sum of $a_{i}^{j}\mu_{i}^{j}$ taken over $j$ such that $a_{i}^{j}>0$, and let $\mu_{i}^{-}$ be the sum of $a_{i}^{j}\mu_{i}^{j}$ taken over $j$ such that $a_{i}^{j}<0$. Then $\mu_{i}^{+}-\mu_{i}^{-}\in a_{i}q_{i}\mathbf{Z}$, so by Lemma 5.3, $\mu_{i}^{+}=\mu_{i}^{-}$, or in other words $\sum_{j}a_{i}^{j}[\mu_{i}^{j}]=0$. So by (0–1), $a_{i}^{j}=0$, so $\sum_{i}a_{i}t_{i}=0$. By (0–2) and (0–3), $(t_{i})_{i}$ is linearly independent, so $a_{i}=0$, establishing (0–1). ∎
Theorem 5.5.
Let $\lambda$ be an oriented lamination which is not a foliation with a dense leaf. Then $\mathcal{E}(\mathcal{M}(\lambda))$ is the set of ergodic measures transverse to sublaminations of $\lambda$.
Proof.
We have already seen that every extreme point of $\mathcal{M}(\lambda)$ is an ergodic measure and so we just need to show the converse. By the Morgan–Shelan decomposition, Theorem 2.12, we may reduce to the case that $\lambda$ either has only closed leaves, or has no closed leaves. If $\lambda$ has only closed leaves, then every ergodic sublamination is a single leaf, and in particular cannot be written as a convex combination of any other ergodic sublaminations. Otherwise, $\lambda$ has no closed leaves, so by Lemma 5.4, $\mathcal{E}(\mathcal{M}(\lambda))$ is finite, say $\{\mu_{1},\dots,\mu_{m}\}$. For each $i\neq j$, there exists a saturated Borel set $E_{ij}$ such that $\mu_{i}(E_{ij})=1$ but $\mu_{j}(E_{ij})=0$. In particular, if $\mu$ is not an extreme point and we write $\mu=\sum_{i}d_{i}\mu_{i}$, then there exist $i\neq j$ such that $d_{i},d_{j}>0$. Then $d_{i}\leq\mu(E_{ij})\leq 1-d_{j}$, so $\mu$ is not ergodic. ∎
5.3. The dual flat $\rho^{*}$
Let $b_{1}:=\dim H_{1}(M,\mathbf{R})$ be the first Betti number of $M$, and let $\rho\in\partial B^{*}$ be a costable unit class. We are going to prove Theorem 1.4: $\rho^{*}$ is a polytope, vertices of $\rho^{*}$ have rational direction iff they are represented by closed leaves of $\lambda_{\rho}$, and $\rho^{*}$ has at most $b_{1}-1$ vertices of irrational direction, Corollary 1.5: if $B$ is strictly convex, then every ergodic calibrated lamination is uniquely ergodic, and Theorem 1.6(1): if $S\subset\partial B$ is flat and $\alpha,\beta\in S$, then their intersection product $\alpha\cdot\beta$ vanishes.
Lemma 5.6.
Let $F$ be a calibration and let $\lambda$ be an ergodic, $F$-calibrated, measured oriented lamination. The following are equivalent:
  1. (1)
    $[\lambda]$ has rational direction.
  2. (2)
    $\lambda$ is a closed hypersurface.
Proof.
If $\lambda$ is a closed hypersurface $N$, then $[\lambda]$ is a rescaling of $[N]$, and $[N]$ is the image of the class of $N$ in $H_{d-1}(M,\mathbf{Z})$.
Conversely, assume that $[\lambda]$ has rational direction. Since $\Gamma$ is finitely generated, we may rescale $M$ suitably so that $[\lambda]$ is a representation $\alpha:\Gamma\to\mathbf{Z}$. Since such representations are identified with homotopy classes of maps $M\to\mathbf{S}^{1}$, the Ruelle-Sullivan current $T_{\lambda}$ takes the form $\mathop{}\!\mathrm{d}u$ for some map $u:M\to\mathbf{S}^{1}$. Let $\tilde{u}\in BV_{\mathrm{loc}}(\tilde{M},\mathbf{R})$ be the universal cover of $u$.
Towards contradiction, let $N$ be a leaf of $\lambda$ which is not closed, and let $\tilde{N}\subset\tilde{M}$ be the preimage of $N$. Since $N$ is not closed and $M$ is compact, there exists $x\in N$ such that $N$ accumulates on itself at $x$, in the sense that for every sufficiently small $r>0$, $N\cap B(x,r)$ has infinitely many connected components. Let $\tilde{x}\in\tilde{M}$ be a point in the preimage of $x$. Thus the set $T$ of $t\in\mathbf{R}$ such that $\partial\{\tilde{u}>t\}$ intersects $B(\tilde{x},r/2)$ is infinite.
We claim that there exists $c>0$ such that for any $t\in\mathbf{R}$ such that $\partial\{\tilde{u}>t\}$ intersects $B(\tilde{x},r/2)$,
$\mathrm{vol}(\partial\{\tilde{u}>t\}\cap B(\tilde{x},r))\geq c.$
To see this, let $\tilde{y}\in\partial\{u>t\}\cap B(\tilde{x},r/2)$. Since $\partial\{\tilde{u}>t\}$ is smooth, its density $\theta$ (in the sense of rectifiable sets) at $\tilde{y}$ is the volume of the unit ball of $\mathbf{R}^{d-1}$. By the monotonicity formula for minimal hypersurfaces [Marques, Theorem 7.11], there exists $A\geq 0$ which only depends on $M$ such that for any $\rho>0$,
$\mathrm{vol}(\partial\{\tilde{u}>t\}\cap B(\tilde{y},\rho))\geq e^{-A\rho^{2}}\theta\rho^{d-1}$
and the claim follows by taking $c:=e^{-Ar^{2}/4}\theta$ and $\rho:=r/2$.
The image of $T$ in $\mathbf{S}^{1}$ is a point, so for any $t,s\in T$, either $t=s$ or $|t-s|\geq 1$. We may assume that there is an infinite increasing sequence $(t_{n})$ in $T$. By the coarea formula [Giusti77, Theorem 1.23],
$\int_{B(x,r)}\star|\mathop{}\!\mathrm{d}\tilde{u}|\geq\sum_{n=0}^{\infty}\int_{t_{n}}^{t_{n+1}}\mathrm{vol}(\partial\{u>t\})\mathop{}\!\mathrm{d}t\geq c\sum_{n=0}^{\infty}(t_{n+1}-t_{n})=\infty,$
which is a contradiction, since $\tilde{u}\in BV_{\mathrm{loc}}(\tilde{M},\mathbf{R})$.
So if $[\lambda]$ has rational direction, then every leaf of $\lambda$ is closed. Since $\lambda$ is ergodic, it follows that $\lambda$ is a single closed leaf. ∎
Proof of Theorem 1.4.
By Lemma 5.6, the map $\Pi:\mathcal{M}(\lambda_{\rho})\to\partial B$ which takes a transverse probability measure to its homology class sends closed hypersurfaces to rational vertices, and measures on sublaminations which are not closed hypersurfaces to irrational vertices. By Corollary 1.3, $\Pi$ is surjective, so in order to bound the number of irrational vertices, we need to bound the number $m$ of measures on sublaminations which are not closed hypersurfaces. So by Lemma 5.4, $m\leq\max(1,b_{1}-1)$.
To complete the proof, we must show that $\rho^{*}$ has finitely many rational vertices. If not, there are infinitely many closed leaves $N_{n}$ of $\lambda_{\rho}$ with distinct homology classes $\alpha_{n}\in\mathcal{E}(\rho^{*})$. The infinite sequence $(\alpha_{n})$ is linearly dependent, so $M\setminus\bigcup_{n}N_{n}$ must be disconnected; therefore there can be no leaf of $\lambda_{\rho}$ which is dense in $M$. There is a measure in $\mathcal{M}(\lambda_{\rho})$ which assigns each $N_{n}$ positive weight, so by the Morgan–Shelan decomposition (Theorem 2.12) and pigeonholing, there is $n\neq m$ such that $N_{n},N_{m}$ are in the same parallel family, so $\alpha_{n}=\alpha_{m}$, a contradiction. ∎
Proof of Corollary 1.5.
Suppose that $B$ is strictly convex, and let $(\kappa,\mu)$ be an ergodic lamination which is $F$-calibrated for some calibration $F$. We may assume that $\kappa$ is not a closed hypersurface, since closed hypersurfaces are uniquely ergodic. Let $\alpha$ be the homology class of $(\kappa,\mu)$ and let $\rho$ be the cohomology class of $F$. By Lemma 3.4, $\mu\in\mathcal{M}(\lambda_{\rho})$, so by Corollary 1.3 and strict convexity of $B$, $\rho^{*}=\{\alpha\}$. Therefore, by Lemma 5.4, $\mathcal{M}(\lambda_{\rho})=\{\mu\}$, so $\kappa$ is uniquely ergodic. ∎
Proof of Theorem 1.6(1).
There exists $\rho\in\partial B^{*}$ such that $S\subseteq\rho^{*}$. By Lemma 3.9, there exist measured sublaminations $\kappa_{\alpha},\kappa_{\beta}$ of $\lambda_{\rho}$, of classes $\alpha,\beta$. Let $\mathop{}\!\mathrm{d}u_{\alpha},\mathop{}\!\mathrm{d}u_{\beta}$ be their Ruelle-Sullivan currents, and suppose that $x$ is in the union of their supports. If $N$ denotes the leaf of $\lambda_{\rho}$ containing $x$, then for $\sigma=\alpha,\beta$,
$\mathop{}\!\mathrm{d}u_{\sigma}(x)=\mathbf{n}_{N}^{\flat}(x)\mu_{\sigma}(x)$
where $\mu_{\sigma}$ is given by (2.3). In particular, $\mathop{}\!\mathrm{d}u_{\alpha}|_{\operatorname{supp}\mathop{}\!\mathrm{d}u_{\beta}}$ is a (possibly distributional) scalar field times $\mathop{}\!\mathrm{d}u_{\beta}$, so $\mathop{}\!\mathrm{d}u_{\alpha}\wedge\mathop{}\!\mathrm{d}u_{\beta}=0$, hence $\alpha\cdot\beta=0$. ∎
5.4. Perimeter-minimizing sets
In this section only, $M$ denotes a complete Riemannian manifold of bounded curvature; we do not assume that $d=7$ or $M$ is closed. A Borel set $U\subseteq M$ is perimeter-minimizing if $1_{U}$ is a function of least gradient. In the proof of Theorem 1.6(2), we shall need an estimate on perimeter-minimizing sets, which we now prove. See [Giusti77, Chapter 5] for the proof when $M$ is an open subset of euclidean space.
Lemma 5.7.
There are constants $\delta,c>0$ which only depend on $d$ such that for every $r\in(0,\delta\|\mathrm{Riem}_{M}\|_{C^{0}}^{-1/2}]$ and $x\in M$ such that $\operatorname{dist}(x,\partial M)>r$,
$\mathrm{vol}(U\cap B(x,r))\geq cr^{d}.$ (5.1)
Proof.
If we take $\delta$ small enough, then we can approximate $B(x,r)$ by a euclidean ball so well that, by the euclidean isoperimetric inequality, for every $0<\rho\leq r$,
$\mathrm{vol}(\partial(U\cap B(x,\rho)))\geq\frac{1}{2c_{d}}\mathrm{vol}(U\cap B(x,\rho))^{\frac{d-1}{d}},$
where $c_{d}>0$ is the euclidean isoperimetric constant. We can reason as in the proof of [Giusti77, Proposition 5.14] to see that for almost every $0<\rho<r$,
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\rho}\mathrm{vol}(U\cap B(x,\rho))\geq\frac{1}{2}\mathrm{vol}(\partial(U\cap B(x,\rho)))\geq\frac{1}{4c_{d}}\mathrm{vol}(U\cap B(x,\rho))^{\frac{d-1}{d}}.$
Let $f(\rho):=\mathrm{vol}(U\cap B(x,\rho))^{1/d}$, so that
$f^{\prime}(\rho)=\frac{\mathrm{vol}(U\cap B(x,\rho)^{\frac{d-1}{d}})}{d}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\rho}\mathrm{vol}(U\cap B(x,\rho))\geq\frac{1}{4dc_{d}}.$
Therefore $f(\rho)\geq\rho/(4dc_{d})$, as desired. ∎
Lemma 5.8.
For every unbounded perimeter-minimizing set $U\subseteq M$, $\mathrm{vol}(U)=\infty$.
Proof.
Let $r:=\min(1,\delta\|\mathrm{Riem}_{M}\|_{C^{0}}^{-1/2})$ where $\delta$ is as in the previous lemma. Since $U$ is unbounded, there is an infinite $2r$-separated set $S\subset U$. Then the set $U\cap\bigcup_{x\in S}B(x,r)$ has infinite volume by (5.1). ∎
5.5. Strict convexity and the derived series of $\pi_{1}(M)$
Let $\Gamma:=\pi_{1}(M)$ and let $(\Gamma^{(n)})$ be the derived series of $\Gamma$. We now prove Theorem 1.6(2): if $\Gamma^{(1)}/\Gamma^{(2)}$ is a torsion group, then the stable unit ball $B$ is strictly convex.
If $\hat{M}\to M$ is a Galois covering space, let $\operatorname{Gal}(\hat{M},M)$ be the Galois group of deck transformations of $\hat{M}\to M$, so $\Gamma=\operatorname{Gal}(\tilde{M},M)$. The universal abelian covering space of $M$, $\tilde{M}^{\rm ab}\to M$, is the Galois covering space such that
$\operatorname{Gal}(\tilde{M}^{\rm ab},M)=\frac{\Gamma}{\Gamma^{(1)}}=H_{1}(M,\mathbf{Z}).$ (5.2)
Since $\mathbf{R}$ is abelian, we have a natural isomorphism
$\operatorname{Hom}(\Gamma,\mathbf{R})$ $=\operatorname{Hom}(\Gamma/\Gamma^{(1)},\mathbf{R})$
and every $\alpha$-equivariant function $u$ on $\tilde{M}$ descends to an $\alpha$-equivariant function $u^{\rm ab}$ on $\tilde{M}^{\rm ab}$. Since $\mathbf{R}$ is abelian and torsion-free, and $\Gamma^{(1)}=\pi_{1}(\tilde{M}^{\rm ab})$, $\Gamma^{(1)}/\Gamma^{(2)}$ is a torsion group iff
$H^{1}(\tilde{M}^{\rm ab},\mathbf{R})=\operatorname{Hom}(\Gamma^{(1)},\mathbf{R})=0.$
The next two lemmata appeared in [Auer12], though the proof of Theorem 1.6(2) does not. Since this manuscript is not publicly available, or complete, we reproduce them here with full credit to the original authors.
Lemma 5.9 ([Auer12]).
Let $u$ be an $\alpha$-equivariant function of least gradient on $\tilde{M}$. Then the set $\{u^{\rm ab}>t\}$ is connected.
Proof.
Suppose that $\{u^{\rm ab}>t\}$ is disconnected. Then $\alpha$ is nonzero: if $\alpha=0$, then $u$ descends to a function of least gradient on $M$, which is constant since $M$ is closed, and then $\{u^{\rm ab}>t\}$ is either empty or $M$, a contradiction.
Let $F$ be a fundamental domain of $M$ in $\tilde{M}^{\rm ab}$. Since $u^{\rm ab}\in L^{\infty}_{\mathrm{loc}}$ [Gorny20, Theorem 4.3] and $F$ is compact, there exists $t_{0}\in\mathbf{R}$ such that $u>t_{0}$ on $F$. Using (5.2) to interpret $H_{1}(M,\mathbf{Z})$ as the group of deck transformations of $\tilde{M}^{\rm ab}$, let
$H:=\bigcup_{\begin{subarray}{c}\rho\in H_{1}(M,\mathbf{Z})\\ \langle\alpha,\rho\rangle>t-t_{0}\end{subarray}}\rho(F).$
For every $x\in H$, there exists $\rho\in H_{1}(M,\mathbf{Z})$ and $y\in F$, $x=\rho(y)$, and then
$u^{\rm ab}(x)=u^{\rm ab}(y)+\langle\alpha,\rho\rangle>t_{0}+t-t_{0}=t$
so $H\subseteq\{u^{\rm ab}>t\}$.
Since $H$ is the set of translations of the connected fundamental domain $F$ by a half-space in the deck group, $H$ is connected. But $\{u^{\rm ab}>t\}$ is disconnected, so there must be a connected component $X$ of $\{u^{\rm ab}>t\}$ which is disjoint from $H$. For any $\rho\in H_{1}(M,\mathbf{Z})$ such that $\langle\alpha,\rho\rangle>0$, $\rho$ sends $\{u^{\rm ab}>t\}$ into itself, since for every $x\in\{u^{\rm ab}>t\}$,
$u^{\rm ab}(\rho(x))=u^{\rm ab}(x)+\langle\alpha,\rho\rangle>t+0=t.$
In particular, $\rho$ sends $X$ into a component of $\{u^{\rm ab}>t\}$. Thus there are two cases to consider:
  1. (0–1)
    There exists $\rho\in H_{1}(M,\mathbf{Z})$ such that $\langle\alpha,\rho\rangle>0$, but $\rho(X)\subseteq X$.
  2. (0–2)
    For every $\rho\in H_{1}(M,\mathbf{Z})$ such that $\langle\alpha,\rho\rangle>0$, $\rho(X)$ is a subset of a component of $\{u^{\rm ab}>t\}$ which is not $X$.
In case (0–1), there exists $x\in X$ and $\theta\in H_{1}(M,\mathbf{Z})$ such that $\theta(x)\in F$; then, for $m$ large,
$\langle\alpha,m\rho-\theta\rangle>t-t_{0},$
so $m\rho(x)\in H$. Therefore $m\rho(X)$ meets $H$, so $X$ meets $H$, a contradiction.
In case (0–2), let $\hat{M}$ be the minimal covering space on which $u$ descends to a function $\hat{u}:\hat{M}\to\mathbf{R}$, thus $\operatorname{Gal}(\hat{M},M)=\Gamma/\ker(\alpha)$. Then $\hat{u}$ has least gradient, and $X$ descends to a component $\hat{X}$ of $\{\hat{u}>y\}$. Then the projection $\psi:\hat{M}\to M$ restricts to an injective map $\hat{X}\to M$. Indeed, if $x_{1},x_{2}\in\hat{X}$ and $\psi(x_{1})=\psi(x_{2})$, then there exists $\rho\in H_{1}(M,\mathbf{Z})/\ker(\alpha)$ such that $\psi(x_{1})=x_{2}$. If $\rho$ is nonzero, then after switching the roles of $x_{1},x_{2}$ as necessary, we may assume that $\rho$ is represented by some $\overline{\rho}\in H_{1}(M,\mathbf{Z})$ such that $\langle\alpha,\overline{\rho}\rangle>0$, a contradiction.
By a straightforward generalization of [BOMBIERI1969, Theorem 1], $\hat{X}$ is perimeter-minimizing. If $\hat{X}$ is bounded, then $\partial\hat{X}$ is competing with the empty set and hence is empty, a contradiction; so $\hat{X}$ is unbounded and therefore has infinite volume by Lemma 5.8. But $\psi$ is an isometry, so $\psi_{*}(\hat{X})$ is an infinite-volume subset of the closed manifold $M$, a contradiction. ∎
Lemma 5.10 ([Auer12]).
Let $u$ be an $\alpha$-equivariant function of least gradient on $\tilde{M}$, and let $\mathscr{G}$ be a set of curves in $\tilde{M}^{\rm ab}$ which spans $H_{1}(\tilde{M}^{\rm ab},\mathbf{R})$. If $\partial\{u^{\rm ab}>t\}$ misses every curve in $\mathscr{G}$, then $\partial\{u^{\rm ab}>t\}$ is connected.
Proof.
We reason by contrapositive. Let $N_{1},N_{2}$ be two distinct components of $\partial\{u^{\rm ab}>t\}$. By Lemma 5.9 (and the analogous result for sublevel sets), $\tilde{M}^{\rm ab}\setminus\partial\{u^{\rm ab}>t\}$ has two components $E_{1},E_{2}$. We construct a curve $\gamma$, transverse to $N_{1}$, which starts at a point $x\in N_{1}$, passes through $E_{1}$, crosses $N_{2}$ into $E_{2}$, and then returns to $x$. In particular $\gamma$ meets $N_{1}$ at a single point, so their intersection number $[\gamma]\cdot[N_{1}]=1$ (possibly after reorienting). Therefore $[\gamma]$ is a nontrivial class in $H_{1}(\tilde{M}^{\rm ab},\mathbf{R})$. ∎
Proof of Theorem 1.6(2).
We prove the contrapositive. If $B$ is not strictly convex, then there exists $\rho\in\partial B^{*}$ such that $\rho^{*}$ is not singleton. In particular, there are two distinct extreme points $\alpha,\beta\in\mathcal{E}(\rho^{*})$, and by Corollary 1.3, we can find distinct ergodic measured oriented sublaminations $\kappa_{\alpha},\kappa_{\beta}$ of $\lambda_{\rho}$. Let $u_{\alpha},u_{\beta}$ be primitives of the Ruelle-Sullivan currents on $\tilde{M}$; by equivariance, they drop to functions $u^{\rm ab}_{\alpha},u^{\rm ab}_{\beta}$ on the universal abelian cover $\tilde{M}^{\rm ab}$.
There must exist leaves $N_{\alpha}$ of $\kappa_{\alpha}$, and $N_{\beta}$ of $\kappa_{\beta}$, which are distinct. If this is not true, then both $\kappa_{\alpha},\kappa_{\beta}$ are the same closed hypersurface, and in particular $\alpha=\beta$, a contradiction. In particular, by adding constants to $u_{\alpha}$ and $u_{\beta}$, we may assume that $\partial\{u_{\alpha}>0\}$ and $\partial\{u_{\beta}>0\}$ descend to distinct leaves of the covering lamination $\tilde{\lambda}^{\rm ab}_{\rho}$. As sets, $\partial\{u_{\alpha}>0\}$ and $\partial\{u_{\beta}>0\}$ are boundaries and therefore are closed; they are also disjoint, since they are distinct leaves of the same lamination. Therefore they are separated by open sets.
Since $\star|\mathop{}\!\mathrm{d}u_{\alpha}|$ and $\star|\mathop{}\!\mathrm{d}u_{\beta}|$ are elements of $\mathcal{M}(\lambda_{\rho})$, so is their mean, which can be expressed as $\star|\mathop{}\!\mathrm{d}u|$ where $u:=(u_{\alpha}+u_{\beta})/2$. In particular, $u$ has least gradient, and
$\partial\{u>0\}=\partial\{u_{\alpha}>0\}\cup\partial\{u_{\beta}>0\}$
and since the right hand side is separated by open sets, $\partial\{u>0\}$ is disconnected. So by Lemma 5.10, $H_{1}(\tilde{M}^{\rm ab},\mathbf{R})$ is nonzero. ∎
5.6. The earthquake norm
The picture which seems to be emerging is that the stable norm is highly analogous to the earthquake norm on the cotangent space $T_{\sigma}^{*}\mathscr{T}_{g}$ to the Teichmüller space of a closed hyperbolic surface $(\Sigma_{g},\sigma)$ of genus $g$. The starting point for this observation is the earthquake theorem, [Kerckhoff83], which asserts that one can identify each $\alpha\in T_{\sigma}^{*}\mathscr{T}_{g}$ uniquely with a measured geodesic lamination in $(\Sigma_{g},\sigma)$; the earthquake norm $\|\alpha\|_{\rm eq}$ is the mass of the corresponding measure. It follows from [daskalopoulos2023, Theorem 1.6] that for each geodesic lamination $\lambda$ in $(\Sigma,\sigma)$, the map $\mathcal{M}(\lambda)\to T_{\sigma}^{*}\mathscr{T}_{g}$ is affine. The dual norm to the earthquake norm is the stretch norm, which is given by infinitesimal minimizing Lipschitz maps, just as the costable norm is given by calibrations.
The earthquake norm satisfies an analogue of Theorem 1.6(1). (Theorem 1.6(2) holds vacuously, since $\pi_{1}(\Sigma_{g})$ is nonabelian and free.)
Theorem 5.11 ([huang2024earthquakemetricteichmullerspace, Theorem 6.1]).
Let $\alpha,\beta\in T_{\sigma}^{*}\mathscr{T}_{g}$ satisfy $\|\alpha\|_{\rm eq}=\|\beta\|_{\rm eq}=1$. The following are equivalent:
  1. (1)
    $\alpha,\beta$ are contained in the same maximal flat of the earthquake unit sphere.
  2. (2)
    As measured geodesic laminations, $\alpha,\beta$ do not intersect transversely.
The intersection product on measured geodesic laminations on $(\Sigma_{g},\sigma)$ corresponds to the Weil-Petersson $2$-form; in particular, it is symplectic. So by Theorem 5.11, every earthquake flat $F$ is contained in a $33g-$ 3-dimensional subspace of $T_{\sigma}^{*}\mathscr{T}_{g}$. Therefore the earthquake norm also satisfies an analogue of Theorem 1.4:
Corollary 5.12.
Let $F\subset T_{\sigma}^{*}\mathscr{T}_{g}$ be a maximal flat of the earthquake unit sphere. Then:
  1. (1)
    There exists a geodesic lamination $\lambda$ such that $F=\mathcal{M}(\lambda)$.
  2. (2)
    $F$ is a convex polytope with at most $33g-$ 3 vertices.
Proof.
Let $(\alpha_{i})$ be a dense sequence in $F$. By Theorem 5.11, if we let $\lambda_{i}:=\bigcup_{j\leq i}\operatorname{supp}\alpha_{j}$, then $\lambda_{i}$ is (the support of) a geodesic lamination such that $\lambda_{i}\subseteq\lambda_{i+1}$. Taking the limit (say, in the Vietoris topology), $\lambda_{i}$ converges to a lamination $\lambda$ for which $\alpha_{i}\in\mathcal{M}(\lambda)$. Since $\mathcal{M}(\lambda)$ is compact, $F\subseteq\mathcal{M}(\lambda)$, and since $F$ is maximal, $\mathcal{M}(\lambda)\subseteq F$. By an easy generalization of Theorem 5.5, $\mathcal{E}(F)$ is the set of ergodic measures on $\mathcal{M}(\lambda)$, so it is linearly independent. But $\mathcal{E}(F)$ is contained in a $33g-$ 3-dimensional vector space, so $\operatorname{card}\mathcal{E}(F)\leq 3g-3$. ∎
Conjecture 5.13.
For every vector $v$ in the stretch unit sphere of $T_{\sigma}\mathscr{T}_{g}$ there exists a hyperbolic structure $\tau\in\mathscr{T}_{g}$ such that $v^{*}$ is the canonical lamination maximally stretched by the homotopy class of the identity $(\Sigma_{g},\sigma)\to(\Sigma_{g},\tau)$.
Conjecture 5.13 seems quite likely to hold in view of Corollaries 1.3 and 5.12. A natural attempt to prove Conjecture 5.13 is to show that there is a diffeomorphism $\exp_{\sigma}:T_{\sigma}\mathscr{T}_{g}\to\mathscr{T}_{g}$ which maps every ray emanating from 00 in $T_{\sigma}\mathscr{T}_{g}$ to a geodesic ray in $\mathscr{T}_{g}$ with respect to Thurston’s stretch metric, such that $v^{*}\subseteq\mathcal{M}(\lambda_{\sigma,\exp_{\sigma}(v)})$. We refer to Pan and Wolf, [pan2022raystructuresteichmullerspace], for more discussion of exponential maps for Thurston’s stretch metric.
[…]